### v1

# Convolution goes higher-order: a biologically inspired mechanism empowers image classification

Abstract
We propose a novel approach to image classification inspired by complex nonlinear biological visual processing, whereby classical convolutional neural networks (CNNs) are equipped with learnable higher-order convolutions. Our model incorporates a Volterra-like expansion of the convolution operator, capturing multiplicative interactions akin to those observed in early and advanced stages of biological visual processing. We evaluated this approach on synthetic datasets by measuring sensitivity to testing higher-order correlations and performance in standard benchmarks (MNIST, FashionMNIST, CIFAR10, CIFAR100 and Imagenette). Our architecture outperforms traditional CNN baselines, and achieves optimal performance with expansions up to 3rd/4th order, aligning remarkably well with the distribution of pixel intensities in natural images. Through systematic perturbation analysis, we validate this alignment by isolating the contributions of specific image statistics to model performance, demonstrating how different orders of convolution process distinct aspects of visual information. Furthermore, Representational Similarity Analysis reveals distinct geometries across network layers, indicating qualitatively different modes of visual information processing. Our work bridges neuroscience and deep learning, offering a path towards more effective, biologically inspired computer vision models. It provides insights into visual information processing and lays the groundwork for neural networks that better capture complex visual patterns, particularly in resource-constrained scenarios.
1 Introduction
”The correlation over space translates into a correlation over order.” (Koenderink & van Doorn, 2018)
Convolutional Neural Networks (CNNs) have become the cornerstone of modern computer vision tasks, demonstrating remarkable performance across a wide range of applications (Krizhevsky et al., 2012), (Simonyan & Zisserman, 2014), (He et al., 2016). The success of CNNs is largely attributed to their ability to capture local patterns and hierarchical features in images through their layered structure and weight sharing mechanism (LeCun et al., 2015), (Zeiler & Fergus, 2014).
However, natural images contain complex correlations and higher-order statistics that extend beyond simple linear relationships (Field, 1993), (Olshausen & Field, 1996), (Schwartz & Simoncelli, 2001). These include texture patterns, edge interactions, and intricate geometric structures that are prevalent in real-world visual scenes. Importantly, these correlations exhibit a hierarchical structure, with higher-order correlations becoming increasingly sparse (Koenderink & van Doorn, 2018). Traditional CNNs, particularly those with limited depth, often struggle to effectively exploit these higher-order correlations (Gatys et al., 2016), (Ustyuzhaninov et al., 2022), (Ding et al., 2020). The limitation of standard CNNs in capturing higher-order correlations stems from their reliance on linear convolutions followed by pointwise nonlinearities (Yu & Salzmann, 2017). Although deeper networks can approximate complex functions through the composition of many layers, this approach may be computationally expensive and require large amounts of training data.
To address this limitation, we propose a learnable extension of the classical linear convolution operator to incorporate higher-order terms, akin to a Volterra expansion (Volterra, 1887), (Volterra, 1930). This approach allows for the direct modeling of multiplicative interactions between neighboring pixels, enabling the network to capture complex local structures more effectively, even in shallow architectures.
We design models by incorporating higher-order convolutions, and compare their performance against standard CNN baselines. We evaluated our model on a range of datasets, beginning with synthetic images typically used to assess sensitivity to higher-order correlations. Subsequently, we extended our experiments to compare our model against CNN baselines on widely-used datasets including MNIST, FashionMNIST, CIFAR10, CIFAR100 and Imagenette. Our findings reveal that these higher-order convolutions improve image classification performance, offering a promising direction for advancing deep learning models in computer vision tasks.
Interestingly, our results show optimal performance at the 3rd or 4th order, aligning remarkably with the distribution of pixel intensities in natural images as analyzed by Koenderink & van Doorn (2018) and Tkačik et al. (2010). This alignment suggests that our approach effectively captures the fundamental structure of visual information in the natural world. Through systematic perturbation analysis, we isolate the contributions of specific image statistics to model performance, providing insight into how different orders of convolution process distinct aspects of visual information. Furthermore, we employ Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) to investigate the internal representations learned by our model compared with standard CNNs, revealing distinct geometries that indicate different strategies for encoding and processing visual information.
Our approach addresses the limitations of simple pointwise nonlinearities in capturing local image structure, especially in networks of limited depth or width. In particular, we show that incorporating such computations at the earliest stages of image processing, even within the first layer of a neural network, yields significant benefits. This finding suggests that models benefit from embedding these computations via direct engagement with the superficial structure of images (Koenderink, 1984).
1.1 Biological Inspiration and Related Work
Our work draws inspiration from the sophisticated mechanisms of biological visual systems, where visual information processing extends beyond simple pointwise nonlinearities and first-order convolutions (Fitzgerald & Clark, 2015; Clark et al., 2014; Lettvin et al., 1959; Peterhans & Von der Heydt, 1993; Levitt et al., 1994). These non-pointwise nonlinearities, which consider spatial relationships between multiple input points, enable more effective extraction and integration of spatial information (Krieger et al., 1997; Zetzsche & Barth, 1990; Barth et al., 1993; Koenderink & Richards, 1988).
This complex processing begins in the retina, where certain cells exhibit responses to visual stimuli that cannot be explained by simple pointwise nonlinearities (Lettvin et al., 1959). In the visual cortex, specialized neurons respond to complex visual features like corners or line endings (Hubel & Wiesel, 1965), performing advanced computations that integrate information across their receptive fields. Such multidimensional, non-pointwise operators are fundamental to visual processing, as emphasized by Zetzsche et al. (1993) and Koenderink & Richards (1988). Similar mechanisms appear across species - fruit flies employ non-pointwise computations for motion detection as early as the second dendritic layer (Fitzgerald & Clark, 2015), suggesting these computations are fundamental for efficient visual processing (Gilbert, 2007). Recent research has further demonstrated that single dendrites can solve complex problems like XOR (Xu et al., 2012; Ran et al., 2020), a capability that pointwise nonlinearities struggle to model (Minsky & Papert, 1969).
In computational research, the concept of nonlinear receptive fields has been explored through various approaches. Berkes & Wiskott (2006) analyzed inhomogeneous quadratic forms as receptive fields, while Zoumpourlis et al. (2017) extended this to CNN-based learning. For video classification, Roheda et al. (2024) introduced Volterra Neural Networks using a low-rank approach. Related developments in fine-grained image classification include bilinear CNN models (Lin et al., 2015) and compact bilinear pooling (Gao et al., 2016), with applications extending to computational neuroscience (Ahrens et al., 2008).
Our approach differs fundamentally from previous methods by embedding higher-order operations within the convolutional operator itself, rather than applying them after feature extraction. This design choice enables the network to learn and apply complex nonlinear transformations throughout its entire depth, creating a more flexible architecture that better constitutes an algorithmic implementation of biological visual processing. By integrating these operations directly into the network’s core building blocks, we enable more efficient feature learning while maintaining architectural simplicity.
1.2 Beyond Pointwise Nonlinearities
Traditional convolutional neural networks typically employ pointwise nonlinearities with weighted sums of inputs as arguments. Expanding these nonlinearities as a polynomial series reveals that different orders of the expansion share the same weights, effectively coupling the terms across orders and leading to inter- and extra-order dependencies:
$\sigma\left(\sum_{i=0}^{2}w_{i}x_{i}\right)$ $\approx\alpha_{0}+\alpha_{1}(w_{1}x_{1}+w_{2}x_{2})+\alpha_{2}(w_{1}x_{1}+w_{2}x_{2})^{2}+\ldots$ (1)
$=\alpha_{0}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{x_{1}}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{2}}+\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{{}^{2}x_{1}}^{2}+\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{{}^{2}x_{2}}^{2}+2\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{1}}\color[rgb]{0,0,0}{x_{2}}+\ldots$
This tied-weight issue becomes particularly problematic when the network is not deep enough or wide enough to compensate, as the precise bounds for enough depth or width depend on the universal approximation theorem and are often challenging to determine exactly (Bahri et al., 2020). Consequently, the standard pointwise nonlinearity with linear summation may fail to capture complex relationships in the data, especially in shallower or narrower networks.
To address this limitation, we propose a novel approach that can be viewed as an implementation of non-pointwise nonlinearities. Our method, which will be described in detail in Section 2, can be seen as a learnable generalized 2-D detector model (Zetzsche & Barth, 1990) representing a receptive field that responds to multiplicative interactions between inputs, similar to how some biological neurons process visual information (see Subsection 1.1). This approach offers a solution for capturing complex correlations at the level of a single mechanism, allowing for more flexible and powerful representations even in compact network architectures.
2 Model Architecture: Higher-Order Convolution
To perform more complex computations while preserving the benefits of locality and weight sharing that are inherent in standard convolutions, we propose the concept of higher-order convolution. This approach extends the traditional convolution operator to include higher-order terms, allowing for more sophisticated feature extraction, and resulting in a stackable layer for empowering CNNs.
In order to explain our higher-order convolutional layer, we start by considering an input image patch P with n elements, reshaped as a vector x. Standard convolution can be expressed in terms of linear filtering, relating input x and output y as follows: $y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}$ where $w^{i}_{1},i=1,...,n$ represents the weights of the convolution kernel.
We expand this function to include quadratic, cubic, and higher-order terms. The general form of this expansion is:
$y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}+\sum_{j=1}^{n}\sum_{i=1}^{n}w_{2}^{ij}x_{i}x_{j}+\sum_{k=1}^{n}\sum_{j=1}^{n}\sum_{i=1}^{n}w_{3}^{ijk}x_{i}x_{j}x_{k}+...$ (2)
Extending this to the entire image, we can express the higher-order convolution operation as:
$Y_{l,m}=b$ $+\sum_{i,j}w_{1}^{ij}X_{l+i,m+j}$
$+\sum_{i,j}\sum_{k,h}w_{2}^{ijkh}X_{l+i,m+j}X_{l+k,m+h}$
$+\sum_{i,j}\sum_{k,h}\sum_{p,q}w_{3}^{ijkhpq}X_{l+i,m+j}X_{l+k,m+h}X_{l+p,m+q}$
+$\ldots$+ … (3)
Here, $Y_{l,m}$ represents the output at position $(l,m)$ in the feature map, $X$ is the input image, and $w_{1}$, $w_{2}$, $w_{3}$ are the learnable weights for the first, second, and third-order terms respectively. The sums over $i$, $j$, $k$, $h$, $p$, $q$ run over the dimensions of the convolution kernels for each order.
By symmetry considerations, the total number of parameters in this expansion grows as: $n_{V}=\frac{(n+p)!}{n!p!}$ where $p$ represents the order of the term and $n$ the size of the kernel. To address potential issues with the relative magnitudes of higher-order terms, we implement a normalization strategy. For kernels of order greater than the 1st, we apply a scaling factor $s$ calculated as $s=\frac{1}{\sqrt{n_{V}}}$. The scaling factor helps to balance the contribution of higher-order terms relative to the first-order (linear) terms, promoting stability and preventing higher-order terms from dominating network behavior during training.
Figure 1 A illustrates this concept for binary image, showing how the classical convolution is extended to include higher-order terms for a single patch. In practice, this operation is applied across the entire image (as shown in the equation above), resulting in a feature map for each order of the expansion, which are then summed together before a standard pointwise nonlinearity as ReLU (see for example, Figure 3 B).
Figure 1: Extending classical convolution (A) A toy example of how higher-order convolution is implemented. (Left) We consider a patch located at the bottom-left of the input image, and by flattening it to a vector. (Right) Classical convolution would correspond to the 1st order term, where a set of weights (the kernel) are learnt and then translated onto other patches (weight-sharing). We extend classical convolution by expanding the first term: for the 2nd order, we compute the outer product of the original vector x with itself (thus obtaining a matrix) and the corresponding set of weights; similarly for the 3rd order, we iterate the same operation one more time and learn the associated set of weights. Feature maps of the three convolutions are then summed up before a standard pointwise nonlinearity as, for example, a ReLU function. (B) Cumulative explained variance of principal components for standard CNN model vs Higher-order CNN models with 2nd and 3rd expansion quantitatively confirms the tied-weight issue for classical models.
At this point, we have outlined the construction of a higher-order convolutional layer. This layer can functionally replace a standard convolutional layer in neural network architectures and is compatible with training through backpropagation.
2.1 A Quantitative Perspective on the Tied-Weight Issue
To demonstrate how a Higher-order CNN (HoCNN) addresses the tied-weight issue introduced in Section 1.2, we conducted a systematic analysis of weight independence in the network’s representations. We generated a fixed input image (32 × 32 pixels) containing binary textures exhibiting all possible 1-, 2-, 3-, and 4-point correlations for 2×2 patches, corresponding to fixing the $x_{i}$ terms in Equation 1. We then randomly initialized the same model architecture 10,000 times and analyzed the activations after the convolutional block (Conv/HoConv + BatchNorm + ReLU + Max Pooling) for two architectures: a CNN with 10 kernels of size 2×2 and a HoCNN with 2 kernels of size 2×2 with higher-order expansions.
To quantify the degree of weight independence, we employed Principal Component Analysis of these activations, measuring the number of principal components (PCs) needed to explain 95% of the variance. Our hypothesis was that tied weights would reduce the number of principal components, while greater weight independence would require more components to capture the same variance. The results confirmed our expectations (see Figure 1 B): the CNN required only 0.9% of PCs (87 out of 9610) to explain 95% of the variance, while the HoCNN required 5.3% of PCs (102 out of 1922) with second-order expansion and 8.3% (159 out of 1922) with third-order expansion. While the CNN has more total components due to its higher channel count (5×), additional analyses with matched channel counts yielded similar results.
These findings quantitatively demonstrate that higher-order convolutions introduce more independent weights, leading to a richer representation space with reduced weight coupling. The increasing number of components needed for variance explanation directly shows how our approach mitigates the tied-weight issue inherent in standard CNNs. Additional analyses with various nonlinearities showed consistent results, further supporting these conclusions (see Appendix A.5).
3 Experiments & Results
We evaluated the performance of our model on multiple datasets, including a synthetic one consisting of textures with higher order correlations and four widely-used benchmarks: MNIST, FashionMNIST, CIFAR-10, CIFAR-100 and Imagenette. This section details our experimental setup and findings.
3.1 Synthetic Dataset: Structured Visual Textures
To rigorously test the sensitivity of our model to higher-order correlations, we generated a synthetic dataset of structured visual textures based on the work of Victor & Conte (2012) (see Figure 2). These textures allow precise control over the type and intensity of contained correlations.
To generate the textures, we utilized a custom software library (Piasini et al., 2021) implementing the method developed by Victor & Conte (2012). This process involves sampling from a distribution of binary textures with specific multipoint correlation probabilities while maximizing entropy. The correlation intensity is parameterized by parity counts of white or black pixels within 1-, 2-, 3-, or 4-pixel tiles (termed gliders, Figure 2)). Notably, two-point and three-point gliders can produce multiple distinct multipoint correlations based on their spatial configurations. Two-point correlations can arise from horizontal, vertical, or oblique gliders, while three-point correlations can give rise to L patterns with various orientations. Our synthetic dataset comprises 2000 training images, 1000 validation images, and 2000 testing images. All reported results are based on the testing set.
We framed the problem as an image classification task with 10 different classes corresponding to various N-point correlations (N ranging from 1 to 4). To establish a baseline, we first tested a CNN model consisting of 3 blocks (2 convolutional - 1st conv layer with 10, 2x2 kernels and 2nd conv layer with 2, 8x2 kernels; both followed by batch normalization, ReLu nonlinearity and max pooling and 1 fully connected layer). Our results indicate (see Table 1 and Figure 2) that this basic architecture fails to correctly discriminate among the different 2-point and 3-point correlations. We then repeated the classification task using three different HoCNNs with expansions up to the 2nd , the 3rd , and the 4th order kernels (2 kernels of size 2x2) and an additional convolutional (conv layer with 2 kernels of size 8x2 + batch norm. + ReLU + pooling) and fully connected layer replicating the structure of the baseline CNN. These networks produced increasingly good performance (Table 1 and Figure 2), highlighting the need for our proposed higher-order convolution approach. For completeness, we report the number of parameters of the different architectures in Table 1.
Table 1: Accuracy and parameters of CNN and HoCNN models on Texture classification
Model Accuracy (%) # Params
CNN (baseline) 59.14 492
HoCNN (up to 2nd order) 82.42 293
HoCNN (up to 3rd order) 89.02 488
HoCNN (up to 4th order) 92.32 1259
Figure 2: Multipoint correlations and glider classification (A) Textures generated with N-point gliders (N ranging from 1 to 4), totaling 10 classes when taking into account parity constraints. (B) Confusion matrices for different models, from top left: baseline CNN; top right: Higher-order CNN (HoCNN) with kernels expanded up to the 2nd order; bottom-left: HoCNN with kernels expanded up to the 3rd order; bottom-right: HoCNN with kernels expanded up to the 4th order. Taken together the four confusion matrices show that higher-orders progressively allow our network to properly capture relevant features for image classification.
3.2 Benchmark Datasets
3.2.1 MNIST, FashionMNIST, CIFAR-10, CIFAR-100
To further validate our approach and ensure its generalizability, we conducted extensive experiments on four well-established benchmark datasets: MNIST, FashionMNIST, CIFAR-10, and CIFAR-100.
We focused on these relatively small datasets to allow for comprehensive testing and analysis. To ensure the robustness of our results and account for the inherent variability in neural network training, we conducted multiple independent runs for each experiment. Specifically, we initialized and trained each model with 50 different random seeds. This approach provides a more reliable estimate of model performance by mitigating the effects of random initializations.
We compared our HoCNN with 3rd order kernels (3×3) against a standard CNN baseline with equivalent kernel size. Notably, extending to 4th order kernels showed no significant improvement in test accuracy. Complete architectural and training details are provided in Appendix A.3, with parameter comparisons in Table 2. Test accuracy results, averaged over 50 realizations (Figure 3 and Table 2), demonstrate consistent performance advantages for HoCNN across all datasets, with particularly notable improvements on the more complex CIFAR-10 and CIFAR-100 tasks. These results suggest fundamentally different representational capabilities in HoCNN (further explored in Section 4). Additional experiments with a deeper and more complex CNN architecture, detailed in Appendix A.1, showed inferior performance compared to our simpler HoCNN, highlighting the efficiency of higher-order feature extraction.
Training Details. For all, experiments previously introduced experiments, we used AdamW optimizer (Loshchilov, 2017) with learning rate 0.001, weight decay 5e-4, batch size 64, and cross-entropy loss. Input images were normalized using z-score standardization, without any additional data augmentation.
3.2.2 Imagenette
To investigate the scalability and real-world applicability of our approach, we conducted experiments with deeper architectures and a more challenging dataset. Specifically, we implemented Higher-order Convolution in a ResNet-18 (He et al., 2016) architecture (HoResNet-18) and evaluated it on Imagenette (Howard, 2020), a subsampled dataset from ImageNet containing 10 classes. The HoResNet-18 achieved better test accuracy compared to the standard ResNet-18 (see Table 2), demonstrating a performance improvement of over 1 percentage point on this challenging dataset.
The HoResNet-18 follows a similar structure to the standard ResNet-18, with a key modification in the first stage, where higher-order residual blocks are used. The remaining stages utilize standard residual blocks. This hybrid approach allows for the incorporation of higher-order convolutions while maintaining the overall structure of the ResNet architecture. The parameter count for both models is provided in Table 2, while architectural details are presented in the Appendix A.3.1. Additionally, our HoResNet-18 architecture achieves superior performance compared to a ResNet-18 adapted VOneNet (Dapello et al., 2020), a biologically-inspired model, while utilizing fewer parameters, as detailed in Appendix A.4.
These results provide strong evidence for the efficacy of our higher-order convolution approach, demonstrating successful scaling to deeper architectures and more complex datasets. By consistently outperforming the baseline across diverse benchmarks and multiple random initializations, our model exhibits robust improvements in image classification performance beyond simple tasks.
Training Details. For Imagenette specifically, images were normalized using mean = [0.5, 0.5, 0.5] and standard deviation = [0.5, 0.5, 0.5], in order to align with VOneNet (Dapello et al., 2020) preprocessing. All the other training details remain consistent with Subsection 3.2.1.
Figure 3: (A) Standard CNN architecture for Image Classification Benchmarks, the Conv Block represents the concatenation of Conv + Batch Norm. + ReLU + Pooling layers (B) Higher-Order CNN architecture: the first layer is split in three branches, representing the three orders of the Volterra expansion. After computing them separately, they’re added together (mimicking once again the Volterra series expansion. Similarly, the HoConv Block represents the concatenation of HoConv + Batch Norm. + ReLU + Pooling layers. For both networks, the number of kernels depend on the classification task.
Table 2: Accuracy comparison (with $\pm$ $\sigma$, standard deviation) of CNN and HoCNN models across image classification benchmarks, averaged over 50 seeds.
Dataset Accuracy (%) $\pm$ $\sigma$ Params
CNN HoCNN CNN HoCNN
MNIST 99.13 $\pm$ 0.09 99.30 $\pm$ 0.07 82,706 80,262
FashionMNIST 90.48 $\pm$ 0.32 90.95 $\pm$ 0.35 82,706 80,262
CIFAR-10 69.76 $\pm$ 0.64 72.87 $\pm$ 0.54 82,706 80,262
CIFAR-100 36.92 $\pm$ 0.71 40.42 $\pm$ 0.84 385,006 383,478
Imagenette 88.13 89.30 11,181,642 11,168,382
4 Image Statistics Sensitivity and Neural Representations
To understand how our Higher-order Convolutional layer (HoConv) processes information differently from standard Convolutional (Conv) layers, we conducted two complementary analyses: an investigation of sensitivity to image statistics through perturbations (akin to adversarial attacks) and a study of representational geometry.
To directly quantify HoCNN’s sensitivity to higher-order image statistics, we conducted a systematic perturbation analysis on the CIFAR-10 test dataset using both architectures (50 pretrained models each), while keeping the network trained on the unperturbed images. We generated synthetic textures with varying orders of statistical correlations (1-, 2-, 3-, and 4-point) and interpolated them with test images at different intensity (I) levels (imgperturbed = (1-I)$\times$×imgCIFAR-10 + I×$\times$×imgperturbation , with I $\in[0.05,0.20]$). While HoCNN achieves superior performance on CIFAR-10 test set images (see Table 2), it shows increased vulnerability to higher-order statistical perturbations. The performance degradation becomes progressively more pronounced as we move from lower to higher-order correlations, with the gap widening at higher perturbation intensities (see Figure 4 A and B for I = 0.12). For instance, with third-order perturbations at I = 0.12, HoCNN’s accuracy drops by 78.0% compared to CNN’s 73.7% (complete results in Appendix A.6).
This enhanced sensitivity to statistical perturbations suggests HoCNN’s increased capacity to leverage higher-order patterns. To better understand how this statistical sensitivity translates into improved classification performance, we examined the model’s internal representations using Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) on CIFAR-10, analyzing activations from 100 test images (10 per class) averaged across 50 model realizations. The Conv and HoConv blocks exhibit distinct representational geometries (Figure 4A & B), with the latter showing more pronounced class-specific patterns. Analysis of pairwise distance distributions (Figure 4D) reveals that HoConv representations are significantly more dispersed in the high-dimensional space, indicating its ability to capture more diverse and discriminative features. Detailed analysis of individual order components and layer-wise correlations is presented in Appendix A.7.
Together, these analyses demonstrate that HoCNN’s superior classification performance stems from its ability to capture and leverage higher-order statistical features in natural images, resulting in richer and more discriminative internal representations.
Figure 4: Perturbation Analysis and Neural Representations. (A) An exemplar test image (CIFAR-10) and perturbed examples with different 1, 2, 3, 4-point correlations statistics and common intensity (I = 0.12). (B) Normalized (wrt unperturbed CIFAR-10) accuracy for Convolutional (Blue) vs Higher-order (Orange) networks: performances of Higher-order network are systematically worse for more structured perturbations at fixed intensity (I = 0.12). (C) Representational Dissimilarity Matrix (RDM) for baseline Convolutional (Conv) Block (Conv layer, Batch Norm, ReLU, Pooling). (D) RDM for the Higher-order Convolutional (HoConv) block. (E) Log Ratio between the two RDMs, capturing different representational geometries between the two blocks. (F) An alternative view of the HoConv block effectiveness versus the Conv block, comparing the pairwise distance distribution across entries of each RDM. The orange distribution, representing the HoConv block, is significantly shifted to the right, supporting the enhanced image classification performances observed in Figure 5 and Table 2.
5 Discussion
Our study introduces a novel approach to image classification by extending convolutional neural networks with higher-order learnable convolutions, inspired by complex nonlinear biological visual processing. Our findings indicate that expansion beyond the 4th order is unnecessary, a conclusion supported by both theoretical insights and empirical evidence. Koenderink & van Doorn (2018) analysis of natural image statistics reveals that the quadric, cubic and quartic power dominates approximately 63%, 35%, and 2% of the pixels, respectively. Our experimental results align with this distribution, showing no significant performance gains beyond the 3rd order in benchmark tasks and the 4th in structured visual textures. Through systematic perturbation analysis, we validate this alignment by isolating the contributions of specific image statistics to model performance. While adversarial perturbations are typically employed to assess and improve network robustness (Szegedy, 2013; Goodfellow et al., 2014) or analyze network behavior (Carter et al., 2019; Santurkar et al., 2019), we leverage them to probe our model’s sensitivity to distinct orders of image statistics, demonstrating how different orders of convolution process specific aspects of the input.
Our approach to structured visual textures relates our findings to current mechanistic models in computational neuroscience, particularly those focusing on the Drosophila visual system, which is known to be sensitive and selective to different 3-point correlations (Fitzgerald & Clark, 2015), (Clark et al., 2014). This connection strengthens the relevance of our model and highlights its potential to provide insight into visual processing in both biological and artificial systems.
The consistent outperformance of our Higher-order CNN (HoCNN) across various datasets demonstrates the effectiveness of incorporating these biologically inspired computations. Representational Similarity Analysis (RSA) reveals distinct geometries, providing evidence that our model processes visual information differently from standard CNNs, suggesting different strategies for encoding and processing visual information.
Future research directions include exploring deeper architectures with multiple higher-order convolutional layers, which presents challenges in balancing increasing nonlinearity across layers. To a certain extent, the scaling technique we employed (see Section 2) revealed better results compared to the case when it was not used. While computational complexity represents a current limitation, we envision future improvements through low-rank tensor decomposition methods. These techniques, which have proven successful in transformer architectures (Hu et al., 2021), could maintain the benefits of higher-order interactions while significantly reducing computational overhead.
In conclusion, our higher-order CNN represents a promising direction for more effective and biologically inspired computer vision models. By exploiting visual patterns and relationships more efficiently, even in shallow architectures, our approach opens new possibilities for advancing the field. Future research should focus on understanding the interplay between different orders of nonlinearity, especially for spatiotemporal inputs (i.e. videos), on exploring more complex architectures, and investigating potential applications beyond image classification, such as object detection or semantic segmentation.
Acknowledgments
We thank Jan Koenderink and Andrea Van Doorn for valuable discussions and for their detailed explanations of their work, which has been a source of inspiration for this research. We are grateful to Felix Wichmann for early discussions during the preliminary stages of this work. The code implementation will be made publicly available upon publication.
Appendix A Appendix
A.1 Computational costs versus traditional methods
We analyze the computational requirements of HoCNN versus traditional CNNs. Our HoCNN architecture for MNIST, Fashion MNIST and CIFAR-10 consists of:
  • First layer: $3\times 3$ Volterra kernels (8 channels) with first, second, and third-order interactions
  • Two standard convolutional blocks (16 and 32 channels)
  • Two fully connected layers with dropout ($p=0.25$)
Total parameters: 80,262
In order to provide a comparison with a more complex CNN model, we additionally tested a deeper version of the Baseline CNN model. In terms of architecture the Deep CNN network consist:
  • Eight convolutional blocks ($3\times 3$ kernels) with channel progression: $8\rightarrow 8\rightarrow 16\rightarrow 16\rightarrow 32\rightarrow 32\rightarrow 64\rightarrow 64$
  • Two fully connected layers with dropout ($p=0.25$)
Total parameters: 108,130 (25% more than HoCNN)
Despite fewer parameters, HoCNN achieves superior accuracy (avg 72.35% vs 71.20%).
Computational complexity scaling:
  • Standard convolution: $\mathcal{O}(C_{out}\times C_{in}\times K^{2}\times H\times W)$
  • Second-order kernels: $\mathcal{O}(C_{out}\times C_{in}\times K^{4}\times H\times W)$
  • Third-order kernels: $\mathcal{O}(C_{out}\times C_{in}\times K^{6}\times H\times W)$
Where: $C_{out}$ = output channels, $C_{in}$ = input channels, $K$ = kernel size, $H,W$ = feature map height and width.
A.2 Placement of Higher-Order Layer & Effects of Larger Kernels in CNN
We conducted extensive experiments to investigate both the optimal placement of higher-order operations within the network hierarchy and the effects of increased kernel sizes in standard CNNs. These experiments provide important insights into the architectural choices made in our main study.
A.2.1 Effect of Higher-Order Layer Placement
We tested the introduction of the higher-order block at different depths in the network, specifically at the second and third convolutional layers. Our results reveal a consistent pattern across datasets: performance gradually degrades as the higher-order block is placed deeper in the architecture. For CIFAR-10, accuracy decreases from 71.17 ± 0.51% when the higher-order block is at the second convolutional layer to 70.29 ± 0.49% at the third layer. Similarly, for CIFAR-100, accuracy drops from 37.27 ± 0.80% to 36.09 ± 0.81%.
This performance degradation can be attributed to earlier layers potentially capturing and overfitting to spurious correlations. While standard convolutions theoretically preserve higher-order correlations, the cascade of nonlinearities and pooling operations can introduce and amplify spurious correlations, potentially compromising the network’s ability to learn meaningful representations at deeper layers.
A.2.2 Impact of Kernel Size
We also investigated whether simply increasing kernel sizes in standard CNNs could achieve benefits similar to our higher-order operations. Despite larger kernels theoretically being capable of capturing higher-order correlations (as demonstrated by improved performance on our synthetic texture dataset), experiments with 5×5 and 7×7 kernels showed decreased test accuracy compared to our baseline CNN. For CIFAR-10, accuracy decreased to 68.72 ± 0.63% with 5×5 kernels and 67.51 ± 0.61% with 7×7 kernels. The pattern was similar for CIFAR-100, where accuracy dropped to 36.04 ± 0.76% with 5×5 kernels and 34.01 ± 0.72% with 7×7 kernels.
Our choice of 2×2 kernels for the main experiments was motivated by the generative process of our synthetic textures, which are based on gliders up to 4 pixels arranged in 2×2 patches. The results suggest that while larger kernels might have the capacity to capture higher-order interactions, the linear nature of standard convolutions makes it challenging to effectively learn these patterns during optimization.
A.2.3 Computational Considerations
The placement of higher-order operations in early layers offers computational advantages. In typical convolutional architectures, the number of channels increases with network depth. Introducing higher-order kernels in later layers, where channel count is larger, would substantially increase both parameter count and computational complexity. Our architectural choice of early placement thus provides an effective balance between computational efficiency and performance.
These findings support our design decision to introduce higher-order operations early in the network, where they can directly process raw image statistics before potential distortion by successive processing layers. The results demonstrate that explicitly modeling higher-order interactions through dedicated operations provides a more effective approach for capturing and leveraging complex statistical features compared to simply increasing kernel sizes in standard CNNs.
A.3 Architectures and Training Procedure: MNIST, FashionMNIST, CIFAR-10, CIFAR-100 & Imagenette
Figure 5: Validation accuracy curves for CNN and HoCNN on image classification benchmarks. Learning curves comparing the performance of the baseline CNN and the proposed HoCNN on MNIST, FashionMNIST, CIFAR-10, and CIFAR-100 datasets. The HoCNN consistently outperforms the CNN across all benchmarks, demonstrating the effectiveness of incorporating higher-order interactions in the convolutional layers.
For image classification benchmarks, we employed a baseline CNN and a HoCNN architecture, with varying complexity depending on the dataset. For MNIST, FashionMNIST, and CIFAR-10, we used a simple architecture consisting of three convolutional blocks (Convolutional Layer, Batch Normalization, ReLU nonlinearity, and Max pooling) followed by two fully connected layers interleaved with dropout and ReLU nonlinearity (see Figure 3). For these datasets, we used 9, 18, and 36 kernels respectively for the convolutional layers of the baseline CNN, while the HoCNN used 8, 16, and 32 kernels. This configuration ensured that the baseline CNN had a comparable, slightly higher number of parameters (82,706 vs. 80,262). For CIFAR-100, we increased the kernel counts to 34, 68, and 136 for the baseline CNN, and 32, 64, and 128 for the HoCNN, resulting in 385,006 and 383,478 parameters respectively.
The validation accuracy curves for the CNN and HoCNN on these benchmarks are provided in the Figure 5, further illustrating the consistent performance advantage of the HoCNN across datasets.
Our training strategy utilized the AdamW optimizer with weight decay and learning rate scheduling (ReduceLROnPlateau). We employed Early Stopping with patience for model selection. For MNIST, FashionMNIST, CIFAR-10, and CIFAR-100, we trained 50 model realizations with different random seeds to ensure reliable performance assessment. For Imagenette, we conducted a single comprehensive evaluation run.
A.3.1 Architectures and Parameter Counts: Imagenette
Standard ResNet-18 architecture, totaling 11,181,642 parameters:
Initial Block:
  • Convolutional layer ($7\times 7$ kernels, 64 channels)
  • Batch normalization
  • ReLU activation
  • Max pooling ($3\times 3$)
Four main stages, each containing 2 residual blocks:
  • Stage 1: 64 channels (2 blocks)
  • Stage 2: 128 channels (2 blocks)
  • Stage 3: 256 channels (2 blocks)
  • Stage 4: 512 channels (2 blocks)
Final layers:
  • Global average pooling
  • Fully connected layer to output classes
Higher-Order ResNet-18 (HoResNet-18), totaling 11,168,382 parameters:
Initial Block (similar to ResNet-18):
  • Convolutional layer ($7\times 7$ kernels, 30 channels)
  • Batch normalization
  • ReLU activation
  • Max pooling ($3\times 3$)
Four main stages, with a hybrid approach:
  • Stage 1: 30 channels with higher-order residual blocks (2 blocks)
  • Stage 2: 128 channels with standard residual blocks (2 blocks)
  • Stage 3: 256 channels with standard residual blocks (2 blocks)
  • Stage 4: 512 channels with standard residual blocks (2 blocks)
Final layers (similar to ResNet-18):
  • Global average pooling
  • Fully connected layer to output classes
A.3.2 Training Setup: Imagenette
Data Preprocessing:
  • Training images: Random resized crop to $224\times 224$, random horizontal flip, normalization
  • Test images: Resize to $256\times 256$, center crop to $224\times 224$, normalization
  • Both models use normalization with mean and std of [0.5, 0.5, 0.5] which results in rescaled pixels with mean close to zero ([-0.07, -0.09, -0.15]) and std close to 0.5 ([0.55, 0.54, 0.58])
Training Configuration:
  • Batch size: 64
  • Loss function: Cross-entropy
  • Optimizer: AdamW with learning rate 0.001 and weight decay 5e-4
  • Learning rate scheduling: ReduceLROnPlateau (halves LR after 5 epochs without improvement)
  • Early stopping: Implemented with 12 epochs patience
A.4 Comparison with VOneNet
We compared the performance of our HoResNet-18 architecture with a ResNet-18 adapted VOneNet (Dapello et al., 2020), a biologically-inspired model designed to better capture the properties of the visual cortex. Our HoResNet-18 model achieved superior accuracy on the Imagenette dataset (89.30%) compared to the VOneNet model (88.02%), while utilizing fewer parameters (11,168,382 vs. 14,445,322).
This improved performance aligns with the findings reported by the VOneNet authors, who observed lower ImageNet accuracy compared to standard ResNet architectures, despite achieving higher brain-scores (i.e., higher explained variance of neural activity recorded from specific brain areas in the visual cortex). This trade-off between ImageNet performance and brain-score can be verified on the Brain-Score framework website by sorting models based on their ImageNet top-1 accuracy.
Our results demonstrate that the HoResNet-18 architecture, equipped with higher-order convolutions, can outperform biologically-inspired models like VOneNet on standard image classification tasks while maintaining a more compact parameter footprint.
A.5 Tied-Weights Issue
A depiction of how Higher-order terms go beyond simple non-pointwise nonlinearities is provided in Figure 6 while PCA analysis with different nonlinearities is provided in Figure 7, obtaining similar qualitative results as with ReLU.
Figure 6: Beyond pointwise nonlinearities (A) Classical pointwise nonlinearity: the nonlinear function is applied after summing the input from the previous layer. A Taylor expansion of this non-linearity reveals the tied-weight problem as addressed in Subsection 1.2. (B) Non-pointwise nonlinearity includes the summation of quadratic or higher-order terms. By modulating the interactions between terms we can untie weights in the Taylor expansion: the optimization process is not saddled with discovering an efficient trade-off between good approximation and balance of different terms.
Figure 7: PCA analysis with different nonlinearities: Leaky Relu, Gelu, Sigmoid, Mish
A.6 Sensitivity to Image Statistics
More detailed results are presented in Table 3 and 4 (in bold largest decrease in performance across models).
Table 3: CNN Model accuracy (%) for each perturbation type and for different intensities I, in parentheses the decrease in accuracy with respect to the case without perturbation (CIFAR-10 avg test accuracy = 69.76%).
Perturbation I=0.05 I=0.09 I=0.12 I=0.16 I=0.20
$\gamma$ 67.0 (-4.0%) 59.9 (-14.1%) 50.5 (-27.6%) 41.4 (-40.7%) 34.1 (-51.2%)
$\beta$- 66.5 (-4.7%) 59.7 (-14.5%) 50.4 (-27.7%) 40.2 (-42.3%) 31.2 (-55.2%)
$\beta|$ 57.8 (-17.1%) 41.5 (-40.5%) 30.8 (-55.8%) 24.6 (-64.7%) 21.0 (-69.9%)
[…]
$\beta/$ 56.6 (-18.9%) 32.6 (-53.2%) 19.0 (-72.8%) 13.9 (-80.1%) 12.2 (-82.5%)
$\theta\llcorner$ 54.2 (-22.2%) 30.9 (-55.6%) 18.6 (-73.3%) 13.8 (-80.2%) 11.9 (-82.9%)
$\theta\ulcorner$ 54.4 (-22.0%) 31.2 (-55.3%) 18.7 (-73.1%) 13.9 (-80.1%) 11.9 (-82.9%)
$\theta\urcorner$ 54.3 (-22.1%) 31.0 (-55.6%) 18.8 (-73.1%) 13.9 (-80.0%) 12.0 (-82.7%)
$\theta\lrcorner$ 54.2 (-22.3%) 30.7 (-55.9%) 18.5 (-73.5%) 13.7 (-80.3%) 11.9 (-83.0%)
$\alpha$ 56.1 (-19.6%) 36.0 (-48.4%) 24.0 (-65.7%) 18.3 (-73.7%) 15.8 (-77.3%)
Table 4: HO-CNN Model accuracy (%) for each perturbation type and for different intensities I, in parentheses the decrease in accuracy with respect to the case without perturbation (CIFAR-10 avg test accuracy = 72.87%).
Perturbation I=0.05 I=0.09 I=0.12 I=0.16 I=0.20
$\gamma$ 70.2 (-3.6%) 62.5 (-14.2%) 51.9 (-28.8%) 41.2 (-43.4%) 32.8 (-55.0%)
$\beta$- 69.3 (-4.9%) 61.7 (-15.3%) 52.1 (-28.5%) 41.3 (-43.3%) 31.7 (-56.5%)
$\beta|$ 58.5 (-19.8%) 40.3 (-44.8%) 29.1 (-60.0%) 23.2 (-68.2%) 20.2 (-72.3%)
[…]
$\beta/$ 58.8 (-19.3%) 30.5 (-58.2%) 16.4 (-77.5%) 12.5 (-82.8%) 11.7 (-83.9%)
$\theta\llcorner$ 56.3 (-22.7%) 29.1 (-60.1%) 16.0 (-78.0%) 12.1 (-83.4%) 10.9 (-85.0%)
$\theta\ulcorner$ 56.3 (-22.8%) 29.1 (-60.0%) 16.2 (-77.8%) 12.2 (-83.2%) 11.0 (-84.9%)
$\theta\urcorner$ 56.4 (-22.6%) 29.4 (-59.6%) 16.4 (-77.5%) 12.3 (-83.1%) 11.0 (-84.9%)
$\theta\lrcorner$ 56.3 (-22.8%) 29.1 (-60.1%) 16.1 (-77.9%) 12.2 (-83.3%) 11.0 (-85.0%)
$\alpha$ 58.4 (-19.8%) 35.8 (-50.9%) 22.5 (-69.2%) 17.0 (-76.6%) 14.9 (-79.6%)
A.7 Neural Representations: Representational Similarity Analysis
To gain deeper insight into how our Higher-order Convolutional layer (HoConv) processes information differently from a standard Convolutional (Conv) layer, we employed Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) on the CIFAR-10 dataset. We analyzed activations from 100 test images (10 per class) averaged across 50 model realizations to mitigate the effects of random initialization (Ding et al., 2021). Our findings reveal distinct representational geometries between the two models. In Figure 8 we consider an alternative distance (hellinger) between the two RDMs, based on the consideration of the fact that values are between 0 and 1.
Figure 8: (Left) Representational Dissimilarity Matrix (RDM) for baseline Convolutional (Conv) Block (Conv layer, Batch Norm, ReLU, Pooling). (Middle) RDM for the Higher-order Convolutional (HoConv) block. (Right) Hellinger Distance between the two RDMs, capturing different representational geometries between the two blocks
A.7.1 Order-wise Representational Differences
The 2nd and 3rd order components of the HoConv layer can be analyzed independently and extract different representations, with progressively smaller mean dissimilarity compared to the baseline Conv layer (Figure 11). These order-wise differences contribute to the overall representational differences between CNN and HoCNN.
Our analysis of neural representations revealed distinct patterns across different order components (see Figure 9 and Figure 11). The standard convolutional layer exhibited a unimodal distribution of pairwise distances, indicating a consistent representation of stimuli. In contrast, the 2nd order component of the HoCNN showed a bimodal distribution, suggesting two distinct scales of representation. This bimodality could indicate the model’s ability to capture both fine-grained similarities and broader categorical differences between stimuli.
The 3rd order component demonstrated a wider spread of distances, with a peak at lower dissimilarity values. This pattern suggests that the 3rd order component might be capturing more subtle relationships or features among the stimuli, complementing the representations of the lower-order components.
These findings point to a more nuanced and hierarchical representation in HoCNNs compared to standard CNNs. The multi-order structure appears to enable the model to simultaneously process information at various scales and levels of abstraction. This could potentially explain the enhanced performance of HoCNNs on certain tasks, as they may be better equipped to capture complex, multi-scale patterns in the input data.
However, further investigation is needed to fully understand the underlying mechanisms and implications of these representational differences. Future work could explore how these distinct representations contribute to specific aspects of model performance, and whether they align with known principles of biological visual processing.
A.7.2 Layer-wise Representational Differences
The representational difference between CNN and HoCNN becomes more explicit when all orders are summed together and fed to the batch normalization, ReLU, and max pooling layers (see Figure 3), giving rise to the output of our HoConv block. We quantify these differences using Representational Dissimilarity Matrices (RDMs) on a subset of the testing set (Figure 4). We compare blocks instead of single layers because the standard convolution is linear, and equipping it with a nonlinearity layer makes the comparison fairer. The Conv and HoConv RDMs show clear structural differences (Figure 4A & B), with the HoConv exhibiting more pronounced class-specific patterns.
A.7.3 Correlation between RDMs across Layers
We further analyzed the average correlation between RDMs across layers of the CNN and HoCNN (Figure 10). The results confirm the divergence in representational structure, with correlations decreasing in higher-order layers. This suggests that the HoConv layer progressively captures more diverse and complex features compared to the standard Conv layer.
Figure 9: Representational dissimilarity matrices across layers. (First row) Representational Dissimilarity Matrices for (left) standard Conv layer (1st layer of the CNN), (middle) 1st order HoConv layer, (right) absolute difference between the two RDMs. (Second row), the RDM for standard Conv is not repeated while (in the middle) we show the RDM for the activations extracted after the 2nd order HoConv layer; (right) the absolute difference between the standard Conv and the 2nd order HoConv RDMs, which presents different representational geometries.
Figure 10: Average Correlation between RDMs for matching layers (Baseline CNN vs HoCNN).
Figure 11: Pairwise distance distribution of RDMs (CIFAR-10). An alternative way to represent RDM matrices is to plot the pairwise distance distribution of their entries. (A) We can see that at the 1st order the representations of the standard Conv Layer (here before the nonlinearity) and of the HoConv Layer are practically the same (1st order HoConv Layer is mathematically equivalent to a Conv Layer). (B) Representations for the 2nd order HoConv Layer are different from the standard convolutional layer: by themselves alone they don’t bring better representations in terms of dissimilarity if compared with the standard Conv Layer ones. (C ) Similar considerations for 3rd order HoConv Layer representations. (D) When considering the output of the two blocks indeed (same as Fig 5. D), the situation is different: the distribution corresponding to the HoConv representations is considerably shifted, leading to “more” dissimilar representations on average.
A.7.4 RSA on Texture Classification
We extended the analysis presented in paragraph 4 of the main paper to the four models trained on synthetic texture data. This analysis aims to highlight the contributions of different kernel orders in texture classification. Following a similar methodology, we randomly selected 10 samples for each texture class from the test set. We then extracted activations at specific layers of the models and used these to construct representational dissimilarity matrices (RDMs), see Figure 12 , Figure 13and Figure 14.
These RDMs provide a visual representation of how the models distinguish between different texture classes at various processing stages. By comparing the RDMs across different layers and between the baseline CNN and HoCNN models, we can gain insights into how higher-order kernels influence the representation and discrimination of texture patterns.
The following figures present these RDMs, allowing for a comparative analysis of how information is processed and transformed through the network layers in both the standard CNN and the HoCNN architectures when dealing with synthetic texture data. This analysis complements our findings from natural image datasets and offers a more comprehensive view of the role of higher-order computations in visual processing tasks.
Figure 12: RDM comparison between baseline CNN and HoCNN 2nd order (summed 1st and 2nd order kernel), on the right absolute difference to highlight representational differences.
Figure 13: RDM comparison between baseline CNN and HoCNN 3rd order (summed 1st, 2nd and 3rd order kernel), on the right absolute difference to highlight representational differences.
Figure 14: RDM comparison between baseline CNN and HoCNN 4th order (summed 1st, 2nd, 3rd and 4th order kernel), on the right absolute difference to highlight representational differences.

### v2

@@
 Abstract
-We propose a novel approach to image classification inspired by complex nonlinear biological visual processing, whereby classical convolutional neural networks (CNNs) are equipped with learnable higher-order convolutions.
-Our model incorporates a Volterra-like expansion of the convolution operator, capturing multiplicative interactions akin to those observed in early and advanced stages of biological visual processing.
-We evaluated this approach on synthetic datasets by measuring sensitivity to testing higher-order correlations and performance in standard benchmarks (MNIST, FashionMNIST, CIFAR10, CIFAR100 and Imagenette).
-Our architecture outperforms traditional CNN baselines, and achieves optimal performance with expansions up to 3rd/4th order, aligning remarkably well with the distribution of pixel intensities in natural images.
-Through systematic perturbation analysis, we validate this alignment by isolating the contributions of specific image statistics to model performance, demonstrating how different orders of convolution process distinct aspects of visual information.
-Furthermore, Representational Similarity Analysis reveals distinct geometries across network layers, indicating qualitatively different modes of visual information processing.
-Our work bridges neuroscience and deep learning, offering a path towards more effective, biologically inspired computer vision models.
-It provides insights into visual information processing and lays the groundwork for neural networks that better capture complex visual patterns, particularly in resource-constrained scenarios.
+We propose a novel enhancement to Convolutional Neural Networks (CNNs) by incorporating learnable higher-order convolutions inspired by nonlinear biological visual processing.
+Our model extends the classical convolution operator using a Volterra-like expansion to capture multiplicative interactions observed in biological vision.
+Through extensive evaluation on standard benchmarks and synthetic datasets, we demonstrate that our architecture consistently outperforms traditional CNN baselines, achieving optimal performance with 3rd/4th order expansions.
+Systematic perturbation analysis and Representational Similarity Analysis reveal that different orders of convolution process distinct aspects of visual information, aligning with the statistical properties of natural images.
+This biologically-inspired approach offers both improved performance and deeper insights into visual information processing.
 1 Introduction
-”The correlation over space translates into a correlation over order.” (Koenderink & van Doorn, 2018)
-Convolutional Neural Networks (CNNs) have become the cornerstone of modern computer vision tasks, demonstrating remarkable performance across a wide range of applications (Krizhevsky et al., 2012), (Simonyan & Zisserman, 2014), (He et al., 2016).
-The success of CNNs is largely attributed to their ability to capture local patterns and hierarchical features in images through their layered structure and weight sharing mechanism (LeCun et al., 2015), (Zeiler & Fergus, 2014).
-However, natural images contain complex correlations and higher-order statistics that extend beyond simple linear relationships (Field, 1993), (Olshausen & Field, 1996), (Schwartz & Simoncelli, 2001).
+"The correlation over space translates into a correlation over order." Koenderink and van Doorn [1]
+Convolutional Neural Networks (CNNs) have become the cornerstone of modern computer vision tasks, demonstrating remarkable performance across a wide range of applications [2], [3], [4].
+The success of CNNs is largely attributed to their ability to capture local patterns and hierarchical features in images through their layered structure and weight sharing mechanism [5], [6].
+However, natural images contain complex correlations and higher-order statistics that extend beyond simple linear relationships [7], [8], [9].
 These include texture patterns, edge interactions, and intricate geometric structures that are prevalent in real-world visual scenes.
-Importantly, these correlations exhibit a hierarchical structure, with higher-order correlations becoming increasingly sparse (Koenderink & van Doorn, 2018).
-Traditional CNNs, particularly those with limited depth, often struggle to effectively exploit these higher-order correlations (Gatys et al., 2016), (Ustyuzhaninov et al., 2022), (Ding et al., 2020).
-The limitation of standard CNNs in capturing higher-order correlations stems from their reliance on linear convolutions followed by pointwise nonlinearities (Yu & Salzmann, 2017).
+Importantly, these correlations exhibit a hierarchical structure, with higher-order correlations becoming increasingly sparse [1].
+Traditional CNNs, particularly those with limited depth, often struggle to effectively exploit these higher-order correlations [10], [11], [12].
+The limitation of standard CNNs in capturing higher-order correlations stems from their reliance on linear convolutions followed by pointwise nonlinearities [13].
 Although deeper networks can approximate complex functions through the composition of many layers, this approach may be computationally expensive and require large amounts of training data.
-To address this limitation, we propose a learnable extension of the classical linear convolution operator to incorporate higher-order terms, akin to a Volterra expansion (Volterra, 1887), (Volterra, 1930).
+To address this limitation, we propose a learnable extension of the classical linear convolution operator to incorporate higher-order terms, akin to a Volterra expansion [14], [15].
 This approach allows for the direct modeling of multiplicative interactions between neighboring pixels, enabling the network to capture complex local structures more effectively, even in shallow architectures.
 We design models by incorporating higher-order convolutions, and compare their performance against standard CNN baselines.
-We evaluated our model on a range of datasets, beginning with synthetic images typically used to assess sensitivity to higher-order correlations.
-Subsequently, we extended our experiments to compare our model against CNN baselines on widely-used datasets including MNIST, FashionMNIST, CIFAR10, CIFAR100 and Imagenette.
+We evaluated our models on a range of datasets, beginning with synthetic images typically used to assess sensitivity to higher-order correlations.
+Subsequently, we extended our experiments to compare our model against CNN baselines on widely-used datasets including MNIST, FashionMNIST, CIFAR10, CIFAR100, Imagenette and Imagenet.
 Our findings reveal that these higher-order convolutions improve image classification performance, offering a promising direction for advancing deep learning models in computer vision tasks.
-Interestingly, our results show optimal performance at the 3rd or 4th order, aligning remarkably with the distribution of pixel intensities in natural images as analyzed by Koenderink & van Doorn (2018) and Tkačik et al.
-(2010).
+Interestingly, our results show optimal performance at the 3rd or 4th order, aligning remarkably with the distribution of pixel intensities in natural images as analyzed by Koenderink and van Doorn [1] and Tkačik et al.
+[16].
 This alignment suggests that our approach effectively captures the fundamental structure of visual information in the natural world.
 Through systematic perturbation analysis, we isolate the contributions of specific image statistics to model performance, providing insight into how different orders of convolution process distinct aspects of visual information.
-Furthermore, we employ Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) to investigate the internal representations learned by our model compared with standard CNNs, revealing distinct geometries that indicate different strategies for encoding and processing visual information.
+Furthermore, we employ Representational Similarity Analysis (RSA) [17] to investigate the internal representations learned by our model compared with standard CNNs, revealing distinct geometries that indicate different strategies for encoding and processing visual information.
 Our approach addresses the limitations of simple pointwise nonlinearities in capturing local image structure, especially in networks of limited depth or width.
 In particular, we show that incorporating such computations at the earliest stages of image processing, even within the first layer of a neural network, yields significant benefits.
-This finding suggests that models benefit from embedding these computations via direct engagement with the superficial structure of images (Koenderink, 1984).
+This finding suggests that models benefit from embedding these computations via direct engagement with the superficial structure of images [18].
 1.1 Biological Inspiration and Related Work
-Our work draws inspiration from the sophisticated mechanisms of biological visual systems, where visual information processing extends beyond simple pointwise nonlinearities and first-order convolutions (Fitzgerald & Clark, 2015; Clark et al., 2014; Lettvin et al., 1959; Peterhans & Von der Heydt, 1993; Levitt et al., 1994).
-These non-pointwise nonlinearities, which consider spatial relationships between multiple input points, enable more effective extraction and integration of spatial information (Krieger et al., 1997; Zetzsche & Barth, 1990; Barth et al., 1993; Koenderink & Richards, 1988).
-This complex processing begins in the retina, where certain cells exhibit responses to visual stimuli that cannot be explained by simple pointwise nonlinearities (Lettvin et al., 1959).
-In the visual cortex, specialized neurons respond to complex visual features like corners or line endings (Hubel & Wiesel, 1965), performing advanced computations that integrate information across their receptive fields.
-Such multidimensional, non-pointwise operators are fundamental to visual processing, as emphasized by Zetzsche et al.
-(1993) and Koenderink & Richards (1988).
-Similar mechanisms appear across species - fruit flies employ non-pointwise computations for motion detection as early as the second dendritic layer (Fitzgerald & Clark, 2015), suggesting these computations are fundamental for efficient visual processing (Gilbert, 2007).
-Recent research has further demonstrated that single dendrites can solve complex problems like XOR (Xu et al., 2012; Ran et al., 2020), a capability that pointwise nonlinearities struggle to model (Minsky & Papert, 1969).
+Our work draws inspiration from biological visual systems, where processing extends beyond simple pointwise nonlinearities and first-order convolutions [19, 20, 21, 22, 23].
+These non-pointwise nonlinearities enable more effective extraction and integration of spatial information [24, 25, 26, 27], beginning in the retina [21] and continuing in the visual cortex where specialized neurons respond to complex features like corners or line endings [28].
 In computational research, the concept of nonlinear receptive fields has been explored through various approaches.
-Berkes & Wiskott (2006) analyzed inhomogeneous quadratic forms as receptive fields, while Zoumpourlis et al.
-(2017) extended this to CNN-based learning.
+Berkes and Wiskott [29] analyzed inhomogeneous quadratic forms as receptive fields, while Li et al.
+[30] provided theoretical insights on Volterra-based convolutions in neural networks.
+Zoumpourlis et al.
+[31] implemented Volterra-inspired convolutions primarily in the first layer of a Wide ResNet-28, but failed to demonstrate significant performance improvements and limited their analysis to second-order expansions.
+In contrast, our work extends this foundation by implementing and systematically analyzing expansions up to the 4th order, where we demonstrate that 3rd/4th order expansions yield optimal performance—a finding we uniquely connect to the statistical properties of natural images documented by Koenderink and van Doorn [1].
 For video classification, Roheda et al.
-(2024) introduced Volterra Neural Networks using a low-rank approach.
-Related developments in fine-grained image classification include bilinear CNN models (Lin et al., 2015) and compact bilinear pooling (Gao et al., 2016), with applications extending to computational neuroscience (Ahrens et al., 2008).
-Our approach differs fundamentally from previous methods by embedding higher-order operations within the convolutional operator itself, rather than applying them after feature extraction.
-This design choice enables the network to learn and apply complex nonlinear transformations throughout its entire depth, creating a more flexible architecture that better constitutes an algorithmic implementation of biological visual processing.
-By integrating these operations directly into the network’s core building blocks, we enable more efficient feature learning while maintaining architectural simplicity.
+[32] introduced Volterra Neural Networks using cascaded 2nd-order filters with low-rank approximations for spatiotemporal tasks, creating separate spatial and temporal streams that are later fused.
+Our approach differs fundamentally, as we focus on static image classification with direct implementation of higher-order terms (up to 4th order) within the convolution operation itself, preserving parameter independence across orders for more flexible feature learning.
+While Roheda et al.
+[32] emphasize computational efficiency through approximations, our work provides deeper insights through systematic perturbation analysis and representational similarity analysis that reveals how different orders process distinct aspects of visual information.
+Related developments in fine-grained image classification include bilinear CNN models [33] and compact bilinear pooling [34], with applications extending to computational neuroscience [35].
+Our approach extends and complements these previous higher-order methods by embedding higher-order operations directly within the convolutional operator, creating independent feature maps for each order that are combined before nonlinearities.
+This design choice enables the network to learn and apply complex nonlinear transformations that better constitute an algorithmic implementation of biological visual processing.
+Through our structured perturbation analysis, representation geometry analysis, and controlled texture experiments, we provide novel insights that go beyond performance improvements, demonstrating why and how higher-order processing benefits visual representation.
 1.2 Beyond Pointwise Nonlinearities
@@
 $\sigma\left(\sum_{i=0}^{2}w_{i}x_{i}\right)\approx\alpha_{0}+\alpha_{1}(w_{1}x_{1}+w_{2}x_{2})+\alpha_{2}(w_{1}x_{1}+w_{2}x_{2})^{2}+\ldots$ (1)
-$=\alpha_{0}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{x_{1}}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{2}}+\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{{}^{2}x_{1}}^{2}+\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{{}^{2}x_{2}}^{2}+2\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{1}}\color[rgb]{0,0,0}{x_{2}}+\ldots$
-This tied-weight issue becomes particularly problematic when the network is not deep enough or wide enough to compensate, as the precise bounds for enough depth or width depend on the universal approximation theorem and are often challenging to determine exactly (Bahri et al., 2020).
+$=\alpha_{0}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{x_{1}}+\color[rgb]{1,0,0}{\alpha_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{2}}+\quad\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,0,0}{{}^{2}x_{1}^{2}}+\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{{}^{2}x_{2}^{2}}+2\color[rgb]{1,.5,0}{\alpha_{2}}\color[rgb]{0,0,1}{w_{1}}\color[rgb]{0,1,0}{w_{2}}\color[rgb]{0,0,0}{x_{1}x_{2}}+\ldots$
+This tied-weight issue becomes particularly problematic when the network is not deep enough or wide enough to compensate, as the precise bounds for enough depth or width depend on the universal approximation theorem and are often challenging to determine exactly [36].
 Consequently, the standard pointwise nonlinearity with linear summation may fail to capture complex relationships in the data, especially in shallower or narrower networks.
 To address this limitation, we propose a novel approach that can be viewed as an implementation of non-pointwise nonlinearities.
-Our method, which will be described in detail in Section 2, can be seen as a learnable generalized 2-D detector model (Zetzsche & Barth, 1990) representing a receptive field that responds to multiplicative interactions between inputs, similar to how some biological neurons process visual information (see Subsection 1.1).
+Our method, which will be described in detail in Section 2, can be seen as a learnable generalized 2-D detector model [25] representing a receptive field that responds to multiplicative interactions between inputs, similar to how some biological neurons process visual information (see Subsection 1.1).
 This approach offers a solution for capturing complex correlations at the level of a single mechanism, allowing for more flexible and powerful representations even in compact network architectures.
-2 Model Architecture: Higher-Order Convolution
+2 Higher-Order Convolution
 To perform more complex computations while preserving the benefits of locality and weight sharing that are inherent in standard convolutions, we propose the concept of higher-order convolution.
-This approach extends the traditional convolution operator to include higher-order terms, allowing for more sophisticated feature extraction, and resulting in a stackable layer for empowering CNNs.
+This approach extends the traditional convolution operator to include higher-order terms, allowing for more sophisticated feature extraction.
 In order to explain our higher-order convolutional layer, we start by considering an input image patch P with n elements, reshaped as a vector x.
 Standard convolution can be expressed in terms of linear filtering, relating input x and output y as follows: $y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}$ where $w^{i}_{1},i=1,...,n$ represents the weights of the convolution kernel.
-We expand this function to include quadratic, cubic, and higher-order terms. The general form of this expansion is:
-$y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}+\sum_{j=1}^{n}\sum_{i=1}^{n}w_{2}^{ij}x_{i}x_{j}+\sum_{k=1}^{n}\sum_{j=1}^{n}\sum_{i=1}^{n}w_{3}^{ijk}x_{i}x_{j}x_{k}+...$ (2)
+We expand this function to include quadratic, cubic, and higher-order terms.
+The general form of this expansion is:
+$y(\textbf{x})=b+\sum_{i=1}^{n}w_{1}^{i}x_{i}+\sum_{j=1}^{n}\sum_{i=1}^{n}w_{2}^{ij}x_{i}x_{j}+\sum_{k=1}^{n}\sum_{j=1}^{n}\sum_{i=1}^{n}w_{3}^{ijk}x_{i}x_{j}x_{k}+$...
 Extending this to the entire image, we can express the higher-order convolution operation as:
-$Y_{l,m}=b$ $+\sum_{i,j}w_{1}^{ij}X_{l+i,m+j}$
-$+\sum_{i,j}\sum_{k,h}w_{2}^{ijkh}X_{l+i,m+j}X_{l+k,m+h}$
-$+\sum_{i,j}\sum_{k,h}\sum_{p,q}w_{3}^{ijkhpq}X_{l+i,m+j}X_{l+k,m+h}X_{l+p,m+q}$
-+$\ldots$+ … (3)
+$Y_{l,m}=b+\sum_{i,j}w_{1}^{ij}X_{l+i,m+j}+\sum_{i,j}\sum_{k,h}w_{2}^{ijkh}X_{l+i,m+j}X_{l+k,m+h}$
+$+\sum_{i,j}\sum_{k,h}\sum_{p,q}w_{3}^{ijkhpq}X_{l+i,m+j}X_{l+k,m+h}X_{l+p,m+q}+\ldots$ (2)
 Here, $Y_{l,m}$ represents the output at position $(l,m)$ in the feature map, $X$ is the input image, and $w_{1}$, $w_{2}$, $w_{3}$ are the learnable weights for the first, second, and third-order terms respectively.
@@
 Figure 1 A illustrates this concept for binary image, showing how the classical convolution is extended to include higher-order terms for a single patch.
-In practice, this operation is applied across the entire image (as shown in the equation above), resulting in a feature map for each order of the expansion, which are then summed together before a standard pointwise nonlinearity as ReLU (see for example, Figure 3 B).
-Figure 1: Extending classical convolution (A) A toy example of how higher-order convolution is implemented.
-(Left) We consider a patch located at the bottom-left of the input image, and by flattening it to a vector.
-(Right) Classical convolution would correspond to the 1st order term, where a set of weights (the kernel) are learnt and then translated onto other patches (weight-sharing).
-We extend classical convolution by expanding the first term: for the 2nd order, we compute the outer product of the original vector x with itself (thus obtaining a matrix) and the corresponding set of weights; similarly for the 3rd order, we iterate the same operation one more time and learn the associated set of weights.
-Feature maps of the three convolutions are then summed up before a standard pointwise nonlinearity as, for example, a ReLU function.
+In practice, this operation is applied across the entire image, resulting in a feature map for each order of the expansion, which are then summed together before a standard pointwise nonlinearity as ReLU (see for example, Figure 2 B).
+This layer can functionally replace a standard convolutional layer in neural network architectures and is compatible with training through backpropagation.
+Figure 1: Extending classical convolution (A) Implementation of higher-order convolution: Input patch is flattened to vector x.
+1st order uses classical convolution with weight-sharing.
+2nd and 3rd orders compute outer products - respectively one and two times - of x with itself and dot product with the corresponding weights.
+Feature maps are then summed before ReLU activation.
 (B) Cumulative explained variance of principal components for standard CNN model vs Higher-order CNN models with 2nd and 3rd expansion quantitatively confirms the tied-weight issue for classical models.
-At this point, we have outlined the construction of a higher-order convolutional layer.
-This layer can functionally replace a standard convolutional layer in neural network architectures and is compatible with training through backpropagation.
+Inset: Example of the 32×32 binary texture image used in this analysis, containing all possible 1-, 2-, 3-, and 4-point correlations for 2×2 patches.
 2.1 A Quantitative Perspective on the Tied-Weight Issue
 To demonstrate how a Higher-order CNN (HoCNN) addresses the tied-weight issue introduced in Section 1.2, we conducted a systematic analysis of weight independence in the network’s representations.
-We generated a fixed input image (32 × 32 pixels) containing binary textures exhibiting all possible 1-, 2-, 3-, and 4-point correlations for 2×2 patches, corresponding to fixing the $x_{i}$ terms in Equation 1.
+We generated a fixed input image (32 × 32 pixels) containing binary textures exhibiting all possible 1-, 2-, 3-, and 4-point correlations for 2×2 patches (see inset in Figure 1B), corresponding to fixing the $x_{i}$ terms in Equation 1.
 We then randomly initialized the same model architecture 10,000 times and analyzed the activations after the convolutional block (Conv/HoConv + BatchNorm + ReLU + Max Pooling) for two architectures: a CNN with 10 kernels of size 2×2 and a HoCNN with 2 kernels of size 2×2 with higher-order expansions.
@@
 3 Experiments & Results
-We evaluated the performance of our model on multiple datasets, including a synthetic one consisting of textures with higher order correlations and four widely-used benchmarks: MNIST, FashionMNIST, CIFAR-10, CIFAR-100 and Imagenette.
-This section details our experimental setup and findings.
+We evaluated our model on a synthetic texture dataset with higher-order correlations and multiple standard benchmarks: MNIST, FashionMNIST, CIFAR-10, CIFAR-100, and ImageNet.
+Additional results on Imagenette and fine-grained classification (CUB-200-2011) are provided in Appendix A.6 and A.7.
 3.1 Synthetic Dataset: Structured Visual Textures
-To rigorously test the sensitivity of our model to higher-order correlations, we generated a synthetic dataset of structured visual textures based on the work of Victor & Conte (2012) (see Figure 2).
+To test our model’s sensitivity to higher-order correlations, we generated a synthetic dataset of structured visual textures based on the work of Victor and Conte [37] (see Figure 2).
 These textures allow precise control over the type and intensity of contained correlations.
-To generate the textures, we utilized a custom software library (Piasini et al., 2021) implementing the method developed by Victor & Conte (2012).
+To generate the textures, we utilized a custom software library [38] implementing the method developed by Victor and Conte [37].
 This process involves sampling from a distribution of binary textures with specific multipoint correlation probabilities while maximizing entropy.
 The correlation intensity is parameterized by parity counts of white or black pixels within 1-, 2-, 3-, or 4-pixel tiles (termed gliders, Figure 2)).
-Notably, two-point and three-point gliders can produce multiple distinct multipoint correlations based on their spatial configurations.
-Two-point correlations can arise from horizontal, vertical, or oblique gliders, while three-point correlations can give rise to L patterns with various orientations.
+Notably, two-point and three-point gliders create distinct correlations through different spatial configurations: horizontal, vertical, or oblique for two-point, and L patterns with various orientations for three-point correlations.
 Our synthetic dataset comprises 2000 training images, 1000 validation images, and 2000 testing images.
-All reported results are based on the testing set.
+All reported results are based on the test set.
 We framed the problem as an image classification task with 10 different classes corresponding to various N-point correlations (N ranging from 1 to 4).
@@
 For completeness, we report the number of parameters of the different architectures in Table 1.
-Table 1: Accuracy and parameters of CNN and HoCNN models on Texture classification
+Table 1: CNN and HoCNN performances on Texture classification
 Model Accuracy (%) # Params
@@
 3.2 Benchmark Datasets
-3.2.1 MNIST, FashionMNIST, CIFAR-10, CIFAR-100
+MNIST, FashionMNIST, CIFAR-10, CIFAR-100
 To further validate our approach and ensure its generalizability, we conducted extensive experiments on four well-established benchmark datasets: MNIST, FashionMNIST, CIFAR-10, and CIFAR-100.
@@
 Training Details.
-For all, experiments previously introduced experiments, we used AdamW optimizer (Loshchilov, 2017) with learning rate 0.001, weight decay 5e-4, batch size 64, and cross-entropy loss.
-Input images were normalized using z-score standardization, without any additional data augmentation.
-3.2.2 Imagenette
-To investigate the scalability and real-world applicability of our approach, we conducted experiments with deeper architectures and a more challenging dataset.
-Specifically, we implemented Higher-order Convolution in a ResNet-18 (He et al., 2016) architecture (HoResNet-18) and evaluated it on Imagenette (Howard, 2020), a subsampled dataset from ImageNet containing 10 classes.
-The HoResNet-18 achieved better test accuracy compared to the standard ResNet-18 (see Table 2), demonstrating a performance improvement of over 1 percentage point on this challenging dataset.
-The HoResNet-18 follows a similar structure to the standard ResNet-18, with a key modification in the first stage, where higher-order residual blocks are used.
-The remaining stages utilize standard residual blocks.
-This hybrid approach allows for the incorporation of higher-order convolutions while maintaining the overall structure of the ResNet architecture.
-The parameter count for both models is provided in Table 2, while architectural details are presented in the Appendix A.3.1.
-Additionally, our HoResNet-18 architecture achieves superior performance compared to a ResNet-18 adapted VOneNet (Dapello et al., 2020), a biologically-inspired model, while utilizing fewer parameters, as detailed in Appendix A.4.
-These results provide strong evidence for the efficacy of our higher-order convolution approach, demonstrating successful scaling to deeper architectures and more complex datasets.
-By consistently outperforming the baseline across diverse benchmarks and multiple random initializations, our model exhibits robust improvements in image classification performance beyond simple tasks.
+For all experiments, we used AdamW optimizer [39] with learning rate 0.001, weight decay 5e-4, batch size 64, cross-entropy loss and early stopping.
+Images were normalized using z-score standardization without data augmentation.
+Experiments used an NVIDIA RTX 4080 Ti GPU, with training times of 10-15 minutes for MNIST/FashionMNIST and 20-30 minutes for CIFAR-10/CIFAR-100 (HoCNN requiring  1.5-2× longer).
+Complexity analysis is in Appendix A.2.
+IMAGENET
+To investigate the scalability and real-world applicability of our approach, we implemented Higher-order Convolution in a ResNet-18 [4] architecture (HoResNet-18) and evaluated it on ImageNet [40].
+To ensure statistical robustness, we conducted 5 independent training runs with different random seeds.
+HoResNet-18 achieved 69.47% $\pm$ 0.09% test accuracy compared to ResNet-18’s 68.62% $\pm$ 0.08%, demonstrating a statistically significant improvement of 0.85% (see Table 2).
+Results on Imagenette and fine-grained classification (CUB-200-2011) are reported in Appendix A.6 and A.7.
+HoResNet-18 maintains the standard ResNet-18 structure, but uses higher-order residual blocks in its first stage and standard blocks thereafter.
+This hybrid design preserves ResNet’s architecture while incorporating higher-order convolutions.
+Parameter counts are shown in Table 2, with architectural details in Appendix A.9.1.
+These results provide strong evidence for the efficacy of our higher-order convolution approach, demonstrating successful scaling to deeper architectures and large-scale datasets.
+To further investigate scalability, we conducted extensive experiments with deeper ResNet architectures (ResNet-34 and ResNet-50), confirming that performance improvements persist across model depths (see detailed results in Appendix A.4).
 Training Details.
-For Imagenette specifically, images were normalized using mean = [0.5, 0.5, 0.5] and standard deviation = [0.5, 0.5, 0.5], in order to align with VOneNet (Dapello et al., 2020) preprocessing.
-All the other training details remain consistent with Subsection 3.2.1.
-Figure 3: (A) Standard CNN architecture for Image Classification Benchmarks, the Conv Block represents the concatenation of Conv + Batch Norm.
-+ ReLU + Pooling layers (B) Higher-Order CNN architecture: the first layer is split in three branches, representing the three orders of the Volterra expansion.
-After computing them separately, they’re added together (mimicking once again the Volterra series expansion.
-Similarly, the HoConv Block represents the concatenation of HoConv + Batch Norm.
-+ ReLU + Pooling layers.
-For both networks, the number of kernels depend on the classification task.
-Table 2: Accuracy comparison (with $\pm$ $\sigma$, standard deviation) of CNN and HoCNN models across image classification benchmarks, averaged over 50 seeds.
-Dataset Accuracy (%) $\pm$ $\sigma$ Params
+For ImageNet, we used standard practices: normalization with mean = [0.485, 0.456, 0.406] and standard deviation = [0.229, 0.224, 0.225]; Stochastic Gradient Descent with momentum = 0.9 as optimizer.
+We conducted 5 independent runs with different random seeds to ensure statistical robustness.
+The ImageNet experiments were run on an NVIDIA A100 GPU, requiring approximately 2-3 days per run.
+The detailed analysis of computational requirements and FLOP measurements for different orders of convolution can be found in Appendix A.2.
+Figure 3: (A) Comparison of standard convolution (Conv) and (B) higher-order convolution (HoConv) blocks.
+Table 2: Accuracy (with $\sigma$, standard deviation over multiple initializations) of CNN and HoCNN models across image classification benchmarks.
+MNIST, FashionMNIST, CIFAR-10, CIFAR-100: 50 runs; ImageNet: 5 runs.
+Dataset Accuracy (%) $\pm$ $\sigma$ # Params
 CNN HoCNN CNN HoCNN
@@
 CIFAR-100 36.92 $\pm$ 0.71 40.42 $\pm$ 0.84 385,006 383,478
-Imagenette 88.13 89.30 11,181,642 11,168,382
+Imagenet 68.62 $\pm$ 0.08 69.47 $\pm$ 0.09 11,689,512 11,676,252
 4 Image Statistics Sensitivity and Neural Representations
 To understand how our Higher-order Convolutional layer (HoConv) processes information differently from standard Convolutional (Conv) layers, we conducted two complementary analyses: an investigation of sensitivity to image statistics through perturbations (akin to adversarial attacks) and a study of representational geometry.
-To directly quantify HoCNN’s sensitivity to higher-order image statistics, we conducted a systematic perturbation analysis on the CIFAR-10 test dataset using both architectures (50 pretrained models each), while keeping the network trained on the unperturbed images.
-We generated synthetic textures with varying orders of statistical correlations (1-, 2-, 3-, and 4-point) and interpolated them with test images at different intensity (I) levels (imgperturbed = (1-I)$\times$×imgCIFAR-10 + I×$\times$×imgperturbation , with I $\in[0.05,0.20]$).
-While HoCNN achieves superior performance on CIFAR-10 test set images (see Table 2), it shows increased vulnerability to higher-order statistical perturbations.
-The performance degradation becomes progressively more pronounced as we move from lower to higher-order correlations, with the gap widening at higher perturbation intensities (see Figure 4 A and B for I = 0.12).
-For instance, with third-order perturbations at I = 0.12, HoCNN’s accuracy drops by 78.0% compared to CNN’s 73.7% (complete results in Appendix A.6).
-This enhanced sensitivity to statistical perturbations suggests HoCNN’s increased capacity to leverage higher-order patterns.
-To better understand how this statistical sensitivity translates into improved classification performance, we examined the model’s internal representations using Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) on CIFAR-10, analyzing activations from 100 test images (10 per class) averaged across 50 model realizations.
-The Conv and HoConv blocks exhibit distinct representational geometries (Figure 4A & B), with the latter showing more pronounced class-specific patterns.
-Analysis of pairwise distance distributions (Figure 4D) reveals that HoConv representations are significantly more dispersed in the high-dimensional space, indicating its ability to capture more diverse and discriminative features.
-Detailed analysis of individual order components and layer-wise correlations is presented in Appendix A.7.
+[…]
 Together, these analyses demonstrate that HoCNN’s superior classification performance stems from its ability to capture and leverage higher-order statistical features in natural images, resulting in richer and more discriminative internal representations.
@@
 (B) Normalized (wrt unperturbed CIFAR-10) accuracy for Convolutional (Blue) vs Higher-order (Orange) networks: performances of Higher-order network are systematically worse for more structured perturbations at fixed intensity (I = 0.12).
-(C) Representational Dissimilarity Matrix (RDM) for baseline Convolutional (Conv) Block (Conv layer, Batch Norm, ReLU, Pooling).
-(D) RDM for the Higher-order Convolutional (HoConv) block.
+(C) Representational Dissimilarity Matrix (RDM) for baseline Convolutional Block (Conv layer, Batch Norm, ReLU, Pooling).
+(D) RDM for the Higher-order Convolutional block.
 (E) Log Ratio between the two RDMs, capturing different representational geometries between the two blocks.
-(F) An alternative view of the HoConv block effectiveness versus the Conv block, comparing the pairwise distance distribution across entries of each RDM.
-The orange distribution, representing the HoConv block, is significantly shifted to the right, supporting the enhanced image classification performances observed in Figure 5 and Table 2.
-5 Discussion
+5 Discussion and Limitations
 Our study introduces a novel approach to image classification by extending convolutional neural networks with higher-order learnable convolutions, inspired by complex nonlinear biological visual processing.
 Our findings indicate that expansion beyond the 4th order is unnecessary, a conclusion supported by both theoretical insights and empirical evidence.
-Koenderink & van Doorn (2018) analysis of natural image statistics reveals that the quadric, cubic and quartic power dominates approximately 63%, 35%, and 2% of the pixels, respectively.
+Koenderink and van Doorn [1]’s analysis of natural image statistics reveals that the quadratic, cubic and quartic power dominates approximately 63%, 35%, and 2% of the pixels, respectively.
 Our experimental results align with this distribution, showing no significant performance gains beyond the 3rd order in benchmark tasks and the 4th in structured visual textures.
 Through systematic perturbation analysis, we validate this alignment by isolating the contributions of specific image statistics to model performance.
-While adversarial perturbations are typically employed to assess and improve network robustness (Szegedy, 2013; Goodfellow et al., 2014) or analyze network behavior (Carter et al., 2019; Santurkar et al., 2019), we leverage them to probe our model’s sensitivity to distinct orders of image statistics, demonstrating how different orders of convolution process specific aspects of the input.
-Our approach to structured visual textures relates our findings to current mechanistic models in computational neuroscience, particularly those focusing on the Drosophila visual system, which is known to be sensitive and selective to different 3-point correlations (Fitzgerald & Clark, 2015), (Clark et al., 2014).
+Our approach to structured visual textures connects to mechanistic models in computational neuroscience, particularly those studying the Drosophila visual system’s sensitivity to 3-point correlations [19, 20].
 This connection strengthens the relevance of our model and highlights its potential to provide insight into visual processing in both biological and artificial systems.
-The consistent outperformance of our Higher-order CNN (HoCNN) across various datasets demonstrates the effectiveness of incorporating these biologically inspired computations.
-Representational Similarity Analysis (RSA) reveals distinct geometries, providing evidence that our model processes visual information differently from standard CNNs, suggesting different strategies for encoding and processing visual information.
-Future research directions include exploring deeper architectures with multiple higher-order convolutional layers, which presents challenges in balancing increasing nonlinearity across layers.
-To a certain extent, the scaling technique we employed (see Section 2) revealed better results compared to the case when it was not used.
-While computational complexity represents a current limitation, we envision future improvements through low-rank tensor decomposition methods.
-These techniques, which have proven successful in transformer architectures (Hu et al., 2021), could maintain the benefits of higher-order interactions while significantly reducing computational overhead.
+The consistent outperformance of our HoCNN across various datasets demonstrates the effectiveness of incorporating these biologically inspired computations.
+RSA reveals distinct geometries, providing evidence that our model processes visual information differently from standard CNNs.
+5.1 Limitations
+Despite the demonstrated benefits, our higher-order convolution approach has several limitations worth acknowledging.
+The primary challenge is computational complexity, as detailed in Appendix A.2.
+Our FLOP analysis reveals that 2nd and 3rd order operations require approximately 5× and 18× more computation than standard convolutions, respectively.
+While we mitigate this by strategic placement of higher-order operations in early network layers, this computational overhead represents a significant trade-off for resource-constrained applications.
+The increasing parameter count with higher-order expansions also introduces scalability challenges.
+Although our implementation leverages kernel symmetry to reduce parameters (e.g., reducing the 3rd order parameter count by 4.42×), memory requirements still increase substantially with order.
+This limits practical applications beyond 4th order expansions, though our experiments indicate minimal benefits beyond this point.
+Additionally, our approach shows increased vulnerability to higher-order statistical perturbations (Section 4), suggesting a trade-off between enhanced feature extraction capability and certain forms of robustness.
+We also observed that placement of higher-order layers significantly impacts performance (Appendix A.3), with deeper placement leading to performance degradation.
+Finally, our current implementation does not employ low-rank approximations or other efficiency techniques commonly used in modern architectures, representing an opportunity for future optimization.
+5.2 Future Directions
+An important consideration is how our higher-order convolution approach relates to Vision Transformers [42], which have emerged as powerful alternatives to CNNs.
+While both capture feature interactions, they operate on fundamentally different principles.
+Our higher-order convolutions explicitly model multiplicative interactions between neighboring pixels through polynomial expansions, whereas transformers employ bilinear operations with global attention.
+From a mathematical perspective, attention mechanisms compute pairwise interactions through the query-key product ($QK^{T}$), which are inherently bilinear, while our approach extends to trilinear and higher-order multiplicative terms within local receptive fields.
+These differences suggest that higher-order convolutions and vision transformers may be complementary rather than competing approaches, potentially offering different advantages for capturing local complex patterns versus global relationships.
+A particularly promising direction involves hybrid architectures that leverage both mechanisms: higher-order convolutions could process local complex patterns and higher-order correlations in early layers, while attention mechanisms capture long-range dependencies and global context in deeper layers.
+This combination could exploit the strengths of both approaches—efficient local feature extraction through higher-order polynomial expansions and flexible global reasoning through self-attention.
+Investigating such architectures could reveal whether these mechanisms provide complementary or redundant information, and identify optimal integration strategies for different vision tasks.
+Beyond architectural hybridization, future research directions include exploring deeper architectures with multiple higher-order convolutional layers and investigating efficiency improvements through low-rank tensor decomposition methods.
+These techniques, which have proven successful in transformer architectures [43], could maintain the benefits of higher-order interactions while significantly reducing computational overhead.
+Additionally, extending our approach to other computer vision tasks beyond image classification, such as object detection and dense prediction tasks, represents an important avenue for validating the generalizability of higher-order convolutions.
 In conclusion, our higher-order CNN represents a promising direction for more effective and biologically inspired computer vision models.
 By exploiting visual patterns and relationships more efficiently, even in shallow architectures, our approach opens new possibilities for advancing the field.
-Future research should focus on understanding the interplay between different orders of nonlinearity, especially for spatiotemporal inputs (i.e.
-videos), on exploring more complex architectures, and investigating potential applications beyond image classification, such as object detection or semantic segmentation.
+Future research should focus on understanding the interplay between different orders of nonlinearity, their relationship with attention mechanisms, and investigating applications across diverse vision tasks.
 Acknowledgments
+We would like to thank the reviewers for their valuable feedback and suggestions that helped improve this work.
 We thank Jan Koenderink and Andrea van Doorn for valuable discussions and for their detailed explanations of their work, which has been a source of inspiration for this research.
 We are grateful to Felix Wichmann for early discussions during the preliminary stages of this work.
-The code implementation will be made publicly available upon publication.
-Appendix A Appendix
-A.1 Computational costs versus traditional methods
+We also thank the Sorbonne Center for Artificial Intelligence (SCAI) for providing essential computational resources.
+Simone Azeglio acknowledges this work was done within the framework of the PostGenAI@Paris project with the reference ANR-23-IACL-0007.
+Additionally, his PhD position was funded by the SCAI, through IDEX Sorbonne Université, project reference ANR-11-IDEX-0004.
+Olivier Marre would like to thank the ERC Consolidator grant DEEPRETINA (101045253) and the Agence Nationale de la Recherche (ANR) for financial support through Chaire Industrielle MyopiaMaster (ANR-22-CHIN-0006), ANR-20-CE37-0018-04-Shooting Star, ANR-22-CE37-0033 NUTRIACT, ANR-22-CE37-0016-01 PerBaCo, and ANR-RetNet4EC.
+Peter Neri acknowledges support from IIT (grant IVXX000701), the Agence nationale de la recherche (ANR-10-LABX-0087, ANR-10-IDEX-0001-02, ANR-22-CE93-0013), and CNRS.
+Ulisse Ferrari acknowledges this work was done within the framework of the PostGenAI@Paris project with the reference ANR-23-IACL-0007, and benefited from financial support by the ANR through grants NatNetNoise (ANR-21-CE37-0024) and IHU FOReSIGHT (ANR-18-IAHU-01).
+Our lab is part of the DIM C-BRAINS, funded by the Conseil Régional d’Île-de-France.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The abstract and introduction clearly state the paper’s contributions - extending convolutional neural networks with higher-order convolutions inspired by biological vision.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: The paper includes a dedicated "Limitations" subsection (Section 5.1) that comprehensively discusses computational complexity challenges, parameter scaling issues, vulnerability to statistical perturbations, layer placement constraints, and the lack of low-rank approximations in the current implementation.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14.
+Justification: The paper doesn’t include formal theoretical results requiring proofs.
+It provides mathematical formulations of higher-order convolutions but doesn’t make claims that require formal theorem-proof structures.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: The paper provides detailed descriptions of the model architecture (Section 2), training procedures (Section 3.2), and datasets used.
+Appendices A-F offer comprehensive architectural specifications, hyperparameters, and implementation details sufficient for reproducing the results.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We included a link to a github repository containing the code used for our experiments in the appendix.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: The paper specifies all relevant training details including optimizer, learning rate, weight decay, batch size, normalization techniques, and data preprocessing steps across all experiments in Section 3.2 and Appendix A.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: Table 2 reports means and standard deviations ($\sigma$) over 50 different random initializations for the benchmark experiments.
+The methodology section (3.2) explicitly states this approach was used to ensure robustness and statistical reliability of the results.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39.
+Justification: The paper includes specific information about compute resources in Section 3.2.
+We specify that experiments on MNIST, FashionMNIST, CIFAR-10, CIFAR-100, and Imagenette were conducted using an NVIDIA RTX 4080 Ti GPU, while ImageNet experiments used an NVIDIA A100 GPU.
+We provide concrete training times: 10-15 minutes per model for MNIST/FashionMNIST, 20-30 minutes for CIFAR-10/CIFAR-100 (with HoCNN models taking 1.5-2× longer than CNN counterparts), 4-5 hours for Imagenette models, and 2-3 days for ImageNet models.
+Additionally, Appendix A.1 provides a comprehensive analysis of computational complexity, including detailed FLOP measurements for different orders of convolution.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conforms with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49.
+Justification: The paper presents foundational research on neural network architectures for image classification, without immediate application-specific societal implications, as the results presented here are not SOTA.
+It focuses on a mathematical extension to convolution operations rather than application-oriented technology.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper doesn’t present models or data with high risk for misuse. It focuses on fundamental image classification techniques using standard benchmark datasets.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59.
+Justification: The paper cites the original sources of all datasets and prior work used, including standard benchmarks (MNIST, CIFAR, ImageNet) and the software library for texture generation (Metex).
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification:
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification:
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification:
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [No]
+  79. Justification:
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A Technical Appendices and Supplementary Material
+A.1 Code Availability & Reproducibility
+The code is available at https://github.com/sazio/HigherOrderConv.
+A.2 Computational costs versus traditional methods
 We analyze the computational requirements of HoCNN versus traditional CNNs. Our HoCNN architecture for MNIST, Fashion MNIST and CIFAR-10 consists of:
@@
 Where: $C_{out}$ = output channels, $C_{in}$ = input channels, $K$ = kernel size, $H,W$ = feature map height and width.
-A.2 Placement of Higher-Order Layer & Effects of Larger Kernels in CNN
+A.2.1 FLOPs Measurements
+To provide a concrete assessment of computational requirements, we conducted detailed benchmarking of our higher-order convolution operations compared to standard convolutions.
+For a single 3×3 convolutional layer processing an ImageNet-sized input:
+  • Order 1 (standard convolution): 0.2312 M FLOPs (baseline)
+  • Order 2: 1.1651 M FLOPs (5.04× standard convolution)
+  • Order 3: 4.3051 M FLOPs (18.62× standard convolution)
+For complete networks on ImageNet-scale inputs:
+  • Standard ResNet-18: 1.82 G FLOPs
+  • HoResNet-18 (1st+2nd order): 2.64 G FLOPs (1.45× increase)
+  • HoResNet-18 (1st+2nd+3rd order): 3.00 G FLOPs (1.65× increase)
+These measurements align with our theoretical complexity analysis while highlighting how strategic placement of higher-order operations primarily in early layers (see A.3) keeps the overall computational increase manageable.
+Additionally, our implementation leverages symmetry in the higher-order kernels, significantly reducing the parameter count:
+  • Order 2: $1.80\times$ reduction ($(3\times 3)^{2}=81\to 45$ unique parameters)
+  • Order 3: $4.42\times$ reduction ($(3\times 3)^{3}=729\to 165$ unique parameters)
+A.2.2 FLOPs-Controlled Experiments
+To isolate the impact of our higher-order convolution approach from simply having additional computational capacity, we conducted an experiment comparing our HoCNN with a deep CNN having comparable FLOPs on CIFAR-10:
+Table 3: Comparison of accuracy with similar computational budgets
+Model FLOPs Accuracy (%)
+Deep CNN 7.09M 71.20
+HoCNN 6.81M 72.06
+These results confirm that the performance gains from higher-order convolutions cannot be attributed merely to increased computational capacity.
+Even with slightly fewer FLOPs, our HoCNN achieves a 1.21% accuracy improvement over the deep CNN.
+This substantiates our claim that higher-order convolutions provide a meaningful inductive bias that enables networks to learn more effective representations than those captured by simply scaling standard architectures.
+A.3 Placement of Higher-Order Layer & Effects of Larger Kernels in CNN
 We conducted extensive experiments to investigate both the optimal placement of higher-order operations within the network hierarchy and the effects of increased kernel sizes in standard CNNs.
 These experiments provide important insights into the architectural choices made in our main study.
-A.2.1 Effect of Higher-Order Layer Placement
+A.3.1 Effect of Higher-Order Layer Placement
 We tested the introduction of the higher-order block at different depths in the network, specifically at the second and third convolutional layers.
@@
 While standard convolutions theoretically preserve higher-order correlations, the cascade of nonlinearities and pooling operations can introduce and amplify spurious correlations, potentially compromising the network’s ability to learn meaningful representations at deeper layers.
-A.2.2 Impact of Kernel Size
+A.3.2 Impact of Kernel Size
 We also investigated whether simply increasing kernel sizes in standard CNNs could achieve benefits similar to our higher-order operations.
@@
 The results suggest that while larger kernels might have the capacity to capture higher-order interactions, the linear nature of standard convolutions makes it challenging to effectively learn these patterns during optimization.
-A.2.3 Computational Considerations
+A.3.3 Computational Considerations
 The placement of higher-order operations in early layers offers computational advantages.
@@
 The results demonstrate that explicitly modeling higher-order interactions through dedicated operations provides a more effective approach for capturing and leveraging complex statistical features compared to simply increasing kernel sizes in standard CNNs.
-A.3 Architectures and Training Procedure: MNIST, FashionMNIST, CIFAR-10, CIFAR-100 & Imagenette
+A.4 ResNet Scaling Experiments
+To thoroughly investigate whether the benefits of our higher-order approach extend to deeper architectures, we conducted comprehensive scaling experiments with ResNet-18, ResNet-34, and ResNet-50 architectures.
+While our main experiments focused on ResNet-18, these additional experiments provide valuable insights into how higher-order convolutions perform as model capacity increases.
+The results comparing standard ResNet and HoResNet across different depths are presented in the following table:
+Table 4: Performance comparison of ResNet vs. HoResNet across architectures of increasing depth
+Dataset Standard ResNet Higher-Order ResNet
+R18 R34 R50 HoR18 HoR34 HoR50
+MNIST 99.34 $\pm$ 0.10 99.35 $\pm$ 0.11 99.38 $\pm$ 0.10 99.38 $\pm$ 0.09 99.40 $\pm$ 0.08 99.41 $\pm$ 0.08
+FashionMNIST 90.85 $\pm$ 0.32 91.15 $\pm$ 0.23 91.20 $\pm$ 0.24 91.25 $\pm$ 0.29 91.32 $\pm$ 0.21 91.34 $\pm$ 0.25
+CIFAR-10 81.26 $\pm$ 0.57 83.02 $\pm$ 0.64 84.10 $\pm$ 0.59 83.05 $\pm$ 0.55 84.06 $\pm$ 0.49 85.50 $\pm$ 0.62
+CIFAR-100 60.77 $\pm$ 0.73 58.05 $\pm$ 0.69 56.46 $\pm$ 0.83 61.85 $\pm$ 0.81 58.35 $\pm$ 0.70 56.55 $\pm$ 0.74
+The results clearly demonstrate that the benefits of our higher-order approach extend beyond ResNet18. Key observations include:
+  • Consistent improvements across model capacities: Our higher-order networks consistently outperform their standard counterparts across all tested architectures and datasets.
+This suggests that the higher-order interactions captured by our approach provide fundamental benefits regardless of network depth.
+  • Scaling behavior: While both standard and higher-order networks show diminishing returns with increasing depth on simpler datasets (as expected), our HoResNets maintain their advantage.
+This indicates that the inductive bias provided by higher-order convolutions remains valuable even as models scale up.
+  • Performance on complex datasets: For CIFAR-10, our HoResNet18 achieves accuracy comparable to ResNet34, while HoResNet50 shows the strongest performance overall.
+This demonstrates the complementary benefits of higher-order operations and increased network depth.
+  • Overfitting considerations: We observe some evidence of overfitting as models get larger, particularly for CIFAR-100, where the limited number of samples per class makes training very deep networks challenging.
+This explains why we initially focused on more compact architectures in our main experiments.
+These findings strongly support our claim that higher-order convolutions provide meaningful improvements that complement traditional scaling strategies like increasing network depth.
+A.5 Corruption Robustness Analysis
+To evaluate the practical robustness of our higher-order models to real-world image corruptions, we conducted comprehensive experiments using the CIFAR-10-C and CIFAR-100-C benchmarks.
+These datasets apply 17 common image corruptions at 5 levels of severity to the standard test sets.
+We report results in terms of normalized accuracy (with respect to performance on the uncorrupted test set).
+Table 5: Normalized accuracy (%) on CIFAR-10-C with severity level 1
+Corruption Type Baseline CNN Higher-Order CNN
+brightness 99.4 $\pm$ 9.2 99.6 $\pm$ 5.2
+contrast 96.0 $\pm$ 14.7 96.8 $\pm$ 6.5
+defocus_blur 98.8 $\pm$ 9.2 99.1 $\pm$ 7.0
+elastic_transform 85.3 $\pm$ 13.3 86.1 $\pm$ 8.5
+fog 97.8 $\pm$ 12.9 98.6 $\pm$ 6.8
+frost 91.2 $\pm$ 24.8 92.9 $\pm$ 13.5
+gaussian_noise 76.2 $\pm$ 77.2 77.5 $\pm$ 30.0
+glass_blur 58.2 $\pm$ 86.7 56.6 $\pm$ 59.0
+impulse_noise 88.3 $\pm$ 25.3 89.6 $\pm$ 11.6
+jpeg_compression 96.3 $\pm$ 12.6 97.0 $\pm$ 6.0
+motion_blur 88.6 $\pm$ 11.2 88.4 $\pm$ 15.7
+pixelate 97.6 $\pm$ 11.1 97.2 $\pm$ 5.9
+saturate 99.5 $\pm$ 10.4 99.9 $\pm$ 5.0
+shot_noise 85.8 $\pm$ 50.4 87.5 $\pm$ 18.9
+snow 90.8 $\pm$ 28.1 90.8 $\pm$ 8.8
+spatter 95.0 $\pm$ 12.8 95.5 $\pm$ 3.1
+zoom_blur 81.5 $\pm$ 30.1 82.5 $\pm$ 18.0
+Table 6: Normalized accuracy (%) on CIFAR-100-C with severity level 1
+Corruption Type Baseline CNN Higher-Order CNN
+brightness 99.1 $\pm$ 9.8 99.2 $\pm$ 12.3
+contrast 93.9 $\pm$ 12.6 94.9 $\pm$ 11.5
+defocus_blur 97.9 $\pm$ 11.3 98.2 $\pm$ 12.4
+elastic_transform 75.1 $\pm$ 13.3 75.6 $\pm$ 10.0
+fog 96.3 $\pm$ 10.3 97.0 $\pm$ 11.0
+frost 84.8 $\pm$ 10.3 86.0 $\pm$ 11.8
+gaussian_noise 60.5 $\pm$ 18.6 64.7 $\pm$ 16.6
+glass_blur 34.6 $\pm$ 21.9 33.2 $\pm$ 20.0
+impulse_noise 72.0 $\pm$ 19.8 73.1 $\pm$ 15.6
+jpeg_compression 92.1 $\pm$ 10.9 93.5 $\pm$ 10.8
+motion_blur 80.7 $\pm$ 18.0 81.5 $\pm$ 12.3
+pixelate 93.4 $\pm$ 14.3 94.0 $\pm$ 12.1
+saturate 98.7 $\pm$ 11.1 98.6 $\pm$ 11.6
+shot_noise 74.0 $\pm$ 16.8 78.3 $\pm$ 15.4
+snow 81.0 $\pm$ 13.3 82.3 $\pm$ 12.8
+spatter 88.5 $\pm$ 13.4 89.4 $\pm$ 11.4
+zoom_blur 73.0 $\pm$ 21.5 74.2 $\pm$ 14.5
+Our results reveal a surprising pattern: despite exhibiting increased sensitivity to our controlled higher-order statistical perturbations (Section 4), HoCNN models demonstrate enhanced robustness to common image corruptions:
+  • For CIFAR-10-C, our HoCNN outperforms the baseline on 13 out of 17 corruptions at severity level 1, with similar patterns observed at higher severity levels.
+  • For CIFAR-100-C, HoCNN shows even more pronounced improvements, outperforming the baseline on 15 out of 17 corruptions at severity level 1.
+  • The improvements are particularly notable for corruptions that preserve important image structure (contrast, blur, fog) while maintaining competitive performance on noise-based corruptions.
+These findings demonstrate that while the higher-order networks are more sensitive to specific statistical manipulations that target their core functionality, they generally exhibit improved robustness to common image corruptions encountered in real-world scenarios.
+This suggests that the richer representational capabilities of higher-order convolutions may actually enhance model reliability in practical applications.
+A.6 Imagenette Results
+To investigate the scalability of our approach on a moderately-sized dataset, we evaluated HoResNet-18 on Imagenette [44], a 10-class subset sampled from ImageNet.
+Imagenette provides a useful intermediate benchmark between the smaller CIFAR datasets and the full ImageNet, allowing for faster experimentation while maintaining the visual complexity of ImageNet images.
+HoResNet-18 achieved 89.30% test accuracy compared to ResNet-18’s 88.13%, demonstrating a +1.17% improvement (Table 7).
+Additionally, our HoResNet-18 architecture achieves superior performance compared to a ResNet-18 adapted VOneNet [45], a biologically-inspired model, while utilizing fewer parameters, as detailed in Appendix A.10.
+Table 7: Imagenette classification accuracy
+Model Accuracy (%) # Params
+ResNet-18 88.13 11,181,642
+HoResNet-18 89.30 11,168,382
+Training Details.
+For Imagenette, images were normalized using mean = [0.5, 0.5, 0.5] and standard deviation = [0.5, 0.5, 0.5], in order to align with VOneNet [45] preprocessing.
+All other training details remain consistent with the benchmarks presented in Section 3.2.
+The Imagenette experiments were conducted on an NVIDIA RTX 4080 Ti GPU with training times of approximately 4-5 hours per model.
+Due to computational constraints, we report single-run results for Imagenette rather than the multiple-run analysis performed for other benchmarks.
+While this limits our ability to assess variance, the improvement is consistent with the trends observed across other datasets, particularly the +0.85% improvement on the full ImageNet with statistical significance confirmed across 5 runs (Section 3.2).
+A.7 Fine-Grained Classification: CUB-200-2011
+To further evaluate our approach on tasks requiring fine-grained visual discrimination, we tested HoResNet-18 on the CUB-200-2011 bird species classification dataset [46].
+Fine-grained classification tasks are particularly relevant for assessing higher-order feature extraction, as they require models to capture subtle visual patterns that distinguish between visually similar categories.
+The CUB-200-2011 dataset contains 11,788 images across 200 bird species, with images exhibiting significant intra-class variation in pose, scale, and background.
+HoResNet-18 achieved 72.31% test accuracy compared to ResNet-18’s 68.24%, demonstrating a substantial improvement of +4.07% (Table 8).
+This larger improvement compared to standard ImageNet classification (+0.85%) suggests that higher-order correlations play an increasingly important role in tasks requiring discrimination of subtle visual patterns.
+The results align with our hypothesis that natural images contain higher-order statistical structures that are particularly relevant for fine-grained recognition tasks.
+Table 8: Fine-grained classification accuracy on CUB-200-2011
+Model Accuracy (%) # Params
+ResNet-18 68.24 11,689,512
+HoResNet-18 72.31 11,676,252
+Training Details.
+For CUB-200-2011, we followed standard fine-grained classification protocols.
+Images were resized to 224×224 pixels and normalized using ImageNet statistics (mean = [0.485, 0.456, 0.406], std = [0.229, 0.224, 0.225]).
+The training procedure matched our ImageNet setup: Stochastic Gradient Descent with momentum = 0.9, initial learning rate = 0.1 with cosine annealing, weight decay = 1e-4, and batch size = 128.
+We used standard data augmentation including random crops, horizontal flips, and color jittering.
+Models were trained for 100 epochs with the best model selected based on validation accuracy.
+Experiments were run on an NVIDIA A100 GPU, requiring approximately 12-18 hours per run.
+Due to computational constraints, we report single-run results for CUB-200-2011 rather than the multiple-run analysis performed for ImageNet.
+However, the substantially larger improvement on this fine-grained task (+4.07% vs.
++0.85% on ImageNet) provides strong evidence that higher-order convolutions are particularly beneficial for capturing subtle visual patterns.
+A.8 Fine-Grained Classification: Extended Analysis
+The CUB-200-2011 dataset presents unique challenges for image classification models due to its fine-grained nature.
+Bird species often differ in subtle plumage patterns, beak shapes, body proportions, and coloration—visual features that require capturing complex spatial relationships beyond simple edges and basic textures.
+This makes CUB-200-2011 an ideal testbed for evaluating whether higher-order convolutions can capture the nuanced statistical structures necessary for fine-grained discrimination.
+Our results on CUB-200-2011 show a +4.07% improvement (HoResNet-18: 72.31% vs.
+ResNet-18: 68.24%), which is notably larger than the +0.85% improvement observed on ImageNet.
+This differential improvement is particularly meaningful: while ImageNet contains diverse object categories with substantial inter-class visual differences, CUB-200-2011 requires discriminating between visually similar bird species.
+The larger performance gain on CUB-200-2011 suggests that higher-order correlations become increasingly important as the visual discrimination task becomes more fine-grained.
+The biological inspiration for our approach aligns well with this finding.
+Biological visual systems are known to excel at fine-grained discrimination tasks—for instance, humans can readily distinguish between similar bird species, fabric textures, or facial identities.
+Our incorporation of higher-order correlation detection mechanisms may capture some of these capabilities, explaining the particularly strong performance on fine-grained tasks.
+Implementation Details: We followed standard fine-grained classification protocols for CUB-200-2011.
+Images were resized to 224×224 pixels and normalized using ImageNet statistics (mean = [0.485, 0.456, 0.406], std = [0.229, 0.224, 0.225]).
+The training procedure matched our ImageNet setup: Stochastic Gradient Descent with momentum = 0.9, initial learning rate = 0.1 with cosine annealing, weight decay = 1e-4, and batch size = 128.
+We used standard data augmentation including random crops, horizontal flips, and color jittering.
+Models were trained for 100 epochs with the best model selected based on validation accuracy.
+Error Analysis: Examining the confusion patterns, we observed that HoResNet-18 particularly excelled at distinguishing species with subtle plumage pattern differences.
+The most common errors for both architectures occurred between species with similar body shapes and coloration but different fine-scale patterns—precisely the scenarios where higher-order correlations would be expected to provide the most benefit.
+This error pattern supports our hypothesis that higher-order convolutions enhance the model’s ability to capture the complex spatial relationships that define fine-grained visual categories.
+A.9 Architectures and Training Procedure: MNIST, FashionMNIST, CIFAR-10, CIFAR-100 & Imagenette
 Figure 5: Validation accuracy curves for CNN and HoCNN on image classification benchmarks.
@@
 For Imagenette, we conducted a single comprehensive evaluation run.
-A.3.1 Architectures and Parameter Counts: Imagenette
+A.9.1 Architectures and Parameter Counts: Imagenette
 Standard ResNet-18 architecture, totaling 11,181,642 parameters:
@@
   • Fully connected layer to output classes
-A.3.2 Training Setup: Imagenette
+A.9.2 Training Setup: Imagenette
 Data Preprocessing:
@@
   • Early stopping: Implemented with 12 epochs patience
-A.4 Comparison with VOneNet
-We compared the performance of our HoResNet-18 architecture with a ResNet-18 adapted VOneNet (Dapello et al., 2020), a biologically-inspired model designed to better capture the properties of the visual cortex.
+A.10 Comparison with VOneNet
+We compared the performance of our HoResNet-18 architecture with a ResNet-18 adapted VOneNet [45], a biologically-inspired model designed to better capture the properties of the visual cortex.
 Our HoResNet-18 model achieved superior accuracy on the Imagenette dataset (89.30%) compared to the VOneNet model (88.02%), while utilizing fewer parameters (11,168,382 vs.
@@
 Our results demonstrate that the HoResNet-18 architecture, equipped with higher-order convolutions, can outperform biologically-inspired models like VOneNet on standard image classification tasks while maintaining a more compact parameter footprint.
-A.5 Tied-Weights Issue
+A.11 Tied-Weights Issue
 A depiction of how Higher-order terms go beyond simple non-pointwise nonlinearities is provided in Figure 6 while PCA analysis with different nonlinearities is provided in Figure 7, obtaining similar qualitative results as with ReLU.
@@
 Figure 7: PCA analysis with different nonlinearities: Leaky Relu, Gelu, Sigmoid, Mish
-A.6 Sensitivity to Image Statistics
-More detailed results are presented in Table 3 and 4 (in bold largest decrease in performance across models).
+A.12 Sensitivity to Image Statistics
+More detailed results are presented in Table 9 and 10 (in bold largest decrease in performance across models).
 Table 9: CNN Model accuracy (%) for each perturbation type and for different intensities I, in parentheses the decrease in accuracy with respect to the case without perturbation (CIFAR-10 avg test accuracy = 69.76%).
@@
 $\beta|$ 57.8 (-17.1%) 41.5 (-40.5%) 30.8 (-55.8%) 24.6 (-64.7%) 21.0 (-69.9%)
-[…]
+$\beta\backslash$ 55.5 (-20.5%) 31.2 (-55.2%) 18.0 (-74.2%) 13.2 (-81.0%) 11.8 (-83.1%)
 $\beta/$ 56.6 (-18.9%) 32.6 (-53.2%) 19.0 (-72.8%) 13.9 (-80.1%) 12.2 (-82.5%)
@@
 $\beta|$ 58.5 (-19.8%) 40.3 (-44.8%) 29.1 (-60.0%) 23.2 (-68.2%) 20.2 (-72.3%)
-[…]
+$\beta\backslash$ 58.5 (-19.7%) 30.6 (-58.0%) 16.5 (-77.4%) 12.6 (-82.7%) 11.7 (-84.0%)
 $\beta/$ 58.8 (-19.3%) 30.5 (-58.2%) 16.4 (-77.5%) 12.5 (-82.8%) 11.7 (-83.9%)
@@
 $\alpha$ 58.4 (-19.8%) 35.8 (-50.9%) 22.5 (-69.2%) 17.0 (-76.6%) 14.9 (-79.6%)
-A.7 Neural Representations: Representational Similarity Analysis
-To gain deeper insight into how our Higher-order Convolutional layer (HoConv) processes information differently from a standard Convolutional (Conv) layer, we employed Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) on the CIFAR-10 dataset.
-We analyzed activations from 100 test images (10 per class) averaged across 50 model realizations to mitigate the effects of random initialization (Ding et al., 2021).
+A.13 Neural Representations: Representational Similarity Analysis
+To gain deeper insight into how our Higher-order Convolutional layer (HoConv) processes information differently from a standard Convolutional (Conv) layer, we employed Representational Similarity Analysis (RSA) [17] on the CIFAR-10 dataset.
+We analyzed activations from 100 test images (10 per class) averaged across 50 model realizations to mitigate the effects of random initialization [47].
 Our findings reveal distinct representational geometries between the two models.
@@
 (Right) Hellinger Distance between the two RDMs, capturing different representational geometries between the two blocks
-A.7.1 Order-wise Representational Differences
+A.13.1 Order-wise Representational Differences
 The 2nd and 3rd order components of the HoConv layer can be analyzed independently and extract different representations, with progressively smaller mean dissimilarity compared to the baseline Conv layer (Figure 11).
@@
 Future work could explore how these distinct representations contribute to specific aspects of model performance, and whether they align with known principles of biological visual processing.
-A.7.2 Layer-wise Representational Differences
+A.13.2 Layer-wise Representational Differences
 The representational difference between CNN and HoCNN becomes more explicit when all orders are summed together and fed to the batch normalization, ReLU, and max pooling layers (see Figure 3), giving rise to the output of our HoConv block.
@@
 The Conv and HoConv RDMs show clear structural differences (Figure 4A & B), with the HoConv exhibiting more pronounced class-specific patterns.
-A.7.3 Correlation between RDMs across Layers
+A.13.3 Correlation between RDMs across Layers
 We further analyzed the average correlation between RDMs across layers of the CNN and HoCNN (Figure 10).
@@
 D), the situation is different: the distribution corresponding to the HoConv representations is considerably shifted, leading to “more” dissimilar representations on average.
-A.7.4 RSA on Texture Classification
+A.13.4 RSA on Texture Classification
 We extended the analysis presented in paragraph 4 of the main paper to the four models trained on synthetic texture data.